### v2

Abstract
We consider linear two-time-scale stochastic approximation algorithms driven by martingale noise. Recent applications in machine learning motivate the need to understand finite-time error rates, but conventional stochastic approximation analysis focus on either asymptotic convergence in distribution or finite-time bounds that are far from optimal. Prior work on asymptotic central limit theorems (CLTs) suggest that two-time-scale algorithms may be able to achieve $1/\sqrt{n}$ error in expectation, with a constant given by the expected norm of the limiting Gaussian vector. However, the best known finite-time rates are much slower. We derive the first non-asymptotic central limit theorem with respect to the Wasserstein-1 distance for two-time-scale stochastic approximation with Polyak-Ruppert averaging. As a corollary, we show that expected error achieved by Polyak-Ruppert averaging decays at rate $1/\sqrt{n}$, which significantly improves on the rates of convergence in prior works.
keywords:
Two-time-scale stochastic approximation, Polyak-Ruppert averaging, Central Limit Theorem
1 Introduction
Stochastic approximation (SA), introduced by Robbins and Monro (1951), is an iterative sample-based approach to find the root (or fixed point) $x^{\star}$ of some unknown operator $F$. In particular, SA iteratively updates $x$, an estimate of $x^{\star}$, by moving along the direction of the sample $F(x;\xi)$ scaled by some step size, where $\xi$ is a random variable representing sampling noise. The simplicity of its implementation has enabled broad applications in many areas including stochastic optimization, machine learning, and reinforcement learning (Sutton and Barto, 2018; Lan, 2020). The convergence properties of SA are well-studied, where the error rate achieved by SA with Polyak-Ruppert (PR) averaging is known to decay faster than that of SA without PR averaging (Polyak and Juditsky, 1992).
In this paper, we study the finite-time error rates of the linear two-time-scale SA (TSA) algorithm, a generalized variant of SA, for solving a system of two coupled linear equations, i.e., we seek to find a pair $(x^{*},y^{*})$ that solves
$\left\{\begin{array}[]{ll}\mathbb{E}_{W}[A_{ff}x^{*}+A_{fs}y^{*}-W]=0,\\ \mathbb{E}_{V}[A_{sf}x^{*}+A_{ss}y^{*}-V]=0,\end{array}\right.$ (3)
where $W,V$ are random vectors with unknown distributions and $A_{ff},A_{fs},A_{sf},A_{ss}$ are the system parameters. In this setting, TSA iteratively updates $(x,y)$ as follows:
$\begin{split}x_{t+1}&=x_{t}-\alpha_{t}\left(A_{ff}x_{t}+A_{fs}y_{t}-W_{t}\right)\\ y_{t+1}&=y_{t}-\gamma_{t}\left(A_{sf}x_{t}+A_{ss}y_{t}-V_{t}\right),\end{split}$ (4)
where $\{(W_{t},V_{t})\}$ is a martingale difference sequence representing the sampling noise. Here, $\alpha_{t}\gg\gamma_{t}$ are two different step sizes. The variable $x$ is often referred to as the fast-time-scale variable (using larger step sizes) while $y$ is called the slow-time-scale variable (using smaller step sizes).
TSA has received a great amount of interest due to its application in many areas where the classic SA is not applicable. Prominent examples include gradient temporal-difference learning (Sutton et al., 2008, 2009; Xu et al., 2019), actor-critic methods in reinforcement learning (Konda and Borkar, 1999; Bhatnagar et al., 2012; Gupta et al., 2019; Wu et al., 2020; Zeng and Doan, 2024b), and gradient descent-ascent methods in min-max optimization for zero-sum games (Zeng et al., 2022). These methods require one iterate being updated using a smaller step size than the other to guarantee convergence.
Unlike the literature on SA where convergence properties are well understood, theoretical results for TSA are not complete. While the benefit of using PR averaging has been established for SA (Polyak and Juditsky, 1992), its benefit has only been demonstrated with respect to asymptotic convergence in distribution for TSA (Mokkadem and Pelletier, 2006). In this paper, we focus on understanding the finite-time error rates at which the PR average of the iterates generated by TSA converge to the desired solution. We use TSA to refer to the sequence $\{(x_{t},y_{t})\}$ in Eq. (4) without Polyak-Ruppert averaging, to distinguish from TSA-PR which is used to refer to the averages $\bar{x}_{n}=n^{-1}\sum_{t=1}^{n}x_{t}$ and $\bar{y}_{n}=n^{-1}\sum_{t=1}^{n}y_{t}$. Our contributions are summarized below.
  1. 1.
    We establish in Theorem 3.2 a finite-time bound on the Wasserstein-1 distance between the scaled deviations $\sqrt{n}(\bar{x}_{n}-x^{*})$ and $\sqrt{n}(\bar{y}_{n}-y^{*})$ generated by TSA-PR and their Gaussian limits. This is the first non-asymptotic central limit theorem (CLT) in the context of two-time-scale algorithms, complementing the asymptotic CLT established in (Mokkadem and Pelletier, 2006).
  2. 2.
    Our bound on the Wasserstein-1 distance implies that the expected error achieved by TSA-PR decays at rate $n^{-1/2}$ (Corollary 3.3), where error is defined to be the norms $\lVert\bar{x}_{n}-x^{*}\rVert$ and $\lVert\bar{y}_{n}-y^{*}\rVert$. This rate significantly improves the rate achieved by TSA without PR averaging, as explained in Section 1.1.1 and Remark 3.4.
  3. 3.
    In proving these results, we also establish in Theorem 3.1 lower and upper bounds on the expected square error (MSE) achieved by TSA (i.e., without PR averaging). When compared to similar bounds in prior works, our result provides a more convenient representation to establish our main results above while using a much simpler proof technique.
1.1 Related Work
In this section, we highlight the differences between our results and prior work. References in the following discussion may assume a model with conditions that differ from ours (Assumptions 13), but nonetheless we refer to their results when they are comparable. The norms $(\lVert x_{n}-x^{*}\rVert,\lVert y_{n}-y^{*}\rVert)$ and $(\lVert\bar{x}_{n}-x^{*}\rVert,\lVert\bar{y}_{n}-y^{*}\rVert)$ are referred to as the errors achieved by TSA and TSA-PR, respectively.
1.1.1 Polyak-Ruppert Averaging in Two-Time-Scale Algorithms
Understanding the optimal rates that can be achieved by a class of algorithms is important for algorithm design. The study of optimal rates in stochastic approximation has been a central focus since (Robbins and Monro, 1951). In the context of single-time-scale algorithms, the question of whether the optimal rate could be achieved was resolved by Polyak and Juditsky (1992), using the PR averaging scheme as proposed by Ruppert (1988); Polyak (1991). Moreover, it was shown that the minimum asymptotic covariance is obtained. The PR averaging scheme was extended to two-time-scale algorithms by Mokkadem and Pelletier (2006), where the authors established asymptotic normality of TSA-PR.
All the above results are asymptotic. The widespread application of stochastic approximation algorithms has spurred interest in understanding finite-time bounds. For single-time-scale algorithms, finite-time bounds on the averaged iterates are known when the step size is constant (Srikant and Ying, 2019; Mou et al., 2020; Durmus et al., 2024). These bounds involve correction terms because constant step sizes are not asymptotically optimal. Finite-time bounds for decaying step sizes was established in (Srikant and Ying, 2019; Chen et al., 2022; Srikant, 2024). Less is known for the more general case of two-time-scale algorithms. Existing finite-time bounds are known only for the TSA without the averaging in Eq. (4); see (Dalal et al., 2018; Kaledin et al., 2020; Haque et al., 2023). Here, we establish the first finite-time bound for the averaged iterates (TSA-PR) generated by two-time-scale algorithms with decreasing step sizes. The rates shown to be achieved by TSA-PR significantly improves that achieved by TSA, akin to the single-time-scale case.
1.1.2 Limit Theorems and Quantitative Bounds
Central limit theorems have been established for both SA and TSA algorithms (Polyak and Juditsky, 1992; Mokkadem and Pelletier, 2006; Hu et al., 2024). But these results are asymptotic, and therefore cannot be applied to rigorously test the significance of repeated trials that halt after a finite number of iterations. Non-asymptotic CLTs were established in (Anastasiou et al., 2019; Samsonov et al., 2024; Srikant, 2024), which capture the normality behavior of single-time-scale algorithms. In this paper (Theorem 3.2), we establish a finite-time bound on the Wasserstein-1 metric for TSA-PR using the martingale CLT in (Srikant, 2024). The reason for considering the Wasserstein-1 distance as in (Srikant, 2024) is that weaker notions of distance considered in (Anastasiou et al., 2019; Samsonov et al., 2024) are not strong enough to deduce explicit bounds such as the expected error. In Corollary 3.3, we show that convergence in the Wasserstein-1 distance can be used to establish both a lower and upper bound on the expected error, capturing its correct magnitude up to exact constants.
2 Model and Preliminaries
In this paper, we consider two time-scale stochastic approximation algorithms of the form
$\begin{split}x_{t+1}&=x_{t}-\alpha_{t}\left(A_{ff}x_{t}+A_{fs}y_{t}-W_{t}\right),\\ y_{t+1}&=y_{t}-\gamma_{t}\left(A_{sf}x_{t}+A_{ss}y_{t}-V_{t}\right).\end{split}$ (5)
While we do not consider the so-called ODE approach here to analyze the system, the assumptions on the system matrices in Eq. (5) are easy to explain by relating the above to a singularly perturbed differential equation (ODE); see (Borkar, 2008):
$\dot{x}_{t}=-(A_{ff}x_{t}+A_{fs}y_{t}),\quad\dot{y}_{t}=-\frac{\gamma}{\alpha}\left(A_{sf}x_{t}+A_{ss}y_{t}\right).$ (6)
In the limit $\gamma/\alpha\to 0$, $x_{t}$ evolves much faster than $y_{t}$. When the system governing $\dot{x}_{t}$ is stable, the slow-time-scale $y_{t}$ is analyzed assuming a stationary solution $x_{\infty}(y)=-A_{ff}^{-1}A_{fs}y$:
$\dot{y}_{t}=-\left(A_{sf}x_{\infty}(y_{t})+A_{ss}y_{t}\right)=-(A_{ss}-A_{sf}A_{ff}^{-1}A_{fs})y_{t}.$ (7)
To ensure that both $x_{t}$ and $y_{t}$ converge to their respective limits, it is therefore assumed that $-A_{ff}$ and $-\Delta=-(A_{ss}-A_{sf}A_{ff}^{-1}A_{fs})$ are both Hürwitz stable, i.e., the eigenvalues lie in the left-half of the complex plane. The rates at which the discretized system in Eq. (5) approach their limits are studied under the same setting, which we now state formally.
Assumption 1 (System Parameters)
The matrix $A_{ff}$ and its Schur complement $\Delta=A_{ss}-A_{sf}A_{ff}^{-1}A_{fs}$ are real and satisfy
$A_{ff}P_{ff}+P_{ff}A_{ff}^{T}=I,\quad\Delta P_{\Delta}+P_{\Delta}\Delta^{T}=I$ (8)
for some positive definite matrices $P_{ff}$ and $P_{\Delta}$. The operator norms of $P_{ff}$ and $P_{\Delta}$ are denoted as $\nu_{ff}$ and $\nu_{\Delta}$, respectively.
The above condition implies a contraction in the weighted norm
$\lVert I-2\alpha A_{ff}\rVert_{P_{ff}}\leq\left(1-\frac{\mu_{ff}}{2}\alpha\right)$ (9)
for some positive constant $\mu_{ff}$ and any sufficiently small $\alpha>0$. The constants $\mu_{ff},\mu_{\Delta}$ are described in more detail in Section A.2.
Assumption 2 (Noise)
Let $\{N_{t}\}\coloneqq\{(W_{t},V_{t})\}_{t=1}^{\infty}$ be a martingale difference sequence drawn independently of the iterates $x_{t}$ and $y_{t}$. We assume that for every $t\geq 1$, $\mathbb{E}\lVert N_{t}\rVert^{2+\beta}<\infty$ for some $\beta\in(1/2,1)$ and that
$\mathbb{E}[W_{t}W_{t}^{T}|\mathcal{H}_{t-1}]=\Gamma_{ff},\mathbb{E}[V_{t}V_{t}^{T}|\mathcal{H}_{t-1}]=\Gamma_{ss},\mathbb{E}[W_{t}V_{t}^{T}|\mathcal{H}_{t-1}]=\Gamma_{fs},$
where $\mathcal{H}_{t}=\{x_{1},y_{1},W_{1},V_{1},\cdots,W_{t},V_{t}\}$ and $\Gamma$ is covariance matrix of $(W_{t},V_{t})$ conditioned on the history.
Our last assumption is a guidance on how to choose the step sizes.
Assumption 3 (Step Size)
The step sizes are chosen to be $\alpha_{t}=\alpha_{1}t^{-a},\gamma_{t}=\gamma_{1}t^{-b}$ for $11/2<a<b<$ 1 and any $\alpha_{1},\gamma_{1}>0$.
Remark 2.1.
Both $x_{n}$ and $y_{n}$ converge even when $b=1$. This parameter is crucial when deducing the $n^{-1}$ rate for the second moment of the slow-time-scale iterate. As we will show that the Polyak-Ruppert averaging scheme achieves this fast rate for both fast- and slow-time-scale variables without having to set $b=1$, we assume $b<1$ which introduces minor technical conditions on the requirement for $\gamma_{1}$.
Here we allow arbitrary $\alpha_{1},\gamma_{1}>0$, but clarify a few technical conditions. Recall the constants $(\mu_{ff},\nu_{ff})$ and $(\mu_{\Delta},\nu_{\Delta})$ associated with $A_{ff}$ and $\Delta$ from Assumption 1. Because $\alpha_{t},\gamma_{t}\to 0$ and $\gamma_{t}/\alpha_{t}\to 0$, there exists a time $t_{0}$ and problem-dependent constants $K_{1},K_{2},K_{3}$ such that the following holds for all $t\geq t_{0}$:
$\alpha_{t}$ $<K_{1}\frac{\mu_{ff}}{\nu_{ff}^{2}},\quad\gamma_{t}<K_{2}\frac{\mu_{\Delta}}{\nu_{\Delta}^{2}},\quad\frac{\gamma_{t}}{\alpha_{t}}=K_{3}\min\left\{\frac{\mu_{ff}}{\nu_{ff}^{2}},\frac{\mu_{\Delta}}{\nu_{\Delta}^{2}}\right\}.$ (10)
Since this condition will be satisfied at some finite-time $t_{0}$, the analysis in this paper holds for all $t\geq t_{0}.$ We set $t_{0}=1$ for simplicity of exposition.
3 Main Results
In this section, we state a finite-time bound on the Wasserstein-1 distance between TSA-PR, scaled appropriately, and its Gaussian limit. We highlight the practical significance of this result by proving lower- and upper-bounds on the expected error. To this end, we first prove that the second moments of TSA error converge to zero at rates determined by the choice of step sizes.
Recall the notations $\hat{x}_{n}=x_{n}-x_{\infty}(y_{n}),\hat{y}_{n}=y_{n}-y^{*}$ to denote the deviation of TSA $(x_{n},y_{n})$ from their targets. Below, we use $\kappa(P)$ to denote the condition number of a matrix $P$.
Theorem 3.1.
Assume 13. Let $\Sigma$ be the asymptotic covariance of $(x_{t}-x^{*},y_{t}-y^{*})$, evaluated as
$A_{ff}\Sigma_{ff}+\Sigma_{ff}A_{ff}^{T}$ $=\Gamma_{ff},$
$A_{ff}\Sigma_{fs}+\Sigma_{ff}A_{sf}^{T}$ $=\Gamma_{fs},$
$\Delta\Sigma_{ss}+\Sigma_{ss}\Delta^{T}+A_{sf}\Sigma_{fs}+\Sigma_{sf}A_{sf}^{T}$ $=\Gamma_{ss}.$ (11)
For some problem-dependent constants $M_{f},M_{s}>0$ and every $n\geq 1$, it holds that
$\begin{split}\lVert\mathbb{E}\hat{x}_{n+1}\hat{x}_{n+1}^{T}-\alpha_{n+1}\Sigma_{ff}\rVert&\leq\kappa(P_{ff})\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert\mathbb{E}\hat{x}_{1}\hat{x}_{1}^{T}-\alpha_{1}\Sigma_{ff}\rVert+M_{f}\gamma_{n},\\ \lVert\mathbb{E}\hat{y}_{n+1}\hat{y}_{n+1}^{T}-\gamma_{n+1}\Sigma_{ss}\rVert&\leq\kappa(P_{\Delta})\prod_{t=1}^{n}\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert\mathbb{E}\hat{y}_{1}\hat{y}_{1}^{T}-\gamma_{1}\Sigma_{ss}\rVert+\frac{M_{s}}{n}\end{split}$ (12)
A proof is provided in Sections B.1B.3. The finite-time bounds in Theorem 3.1 are needed to derive our next result in Theorem 3.2. Specifically, it is implied that
$\lim_{n\to\infty}\lVert\mathbb{E}\hat{x}_{n+1}\hat{x}_{n+1}^{T}-\alpha_{n+1}\Sigma_{ff}\rVert=0,\quad\lim_{n\to\infty}\lVert\mathbb{E}\hat{y}_{n+1}\hat{y}_{n+1}^{T}-\gamma_{n+1}\Sigma_{ss}\rVert=0.$ (13)
Note that the nature of Theorem 2 is similar to Theorem 4.1 in (Haque et al., 2023). However, the latter is not directly applicable to our setting. Our presentation provides a more convenient step to establish the convergence rates in Theorem 3.2. In addition, we develop a simple proof technique for Theorem 3.1 compared to existing proofs for finite-time mean square error rates.
Next, we present a non-asymptotic CLT for TSA-PR. To our knowledge, Mokkadem and Pelletier (2006) are the only authors who analyzed TSA-PR in the context of two-time-scale SA, where the authors prove its asymptotic behavior. Utilizing the error bounds in Theorem 3.1 for TSA, we establish a non-asymptotic CLT on the deviation $\sqrt{n}(\bar{z}_{n}-z^{*})$ from a Gaussian variable $\Sigma^{1/2}Z$ by obtaining a finite-time bound on the Wasserstein-1 distance
$d_{1}\left(\sqrt{n}(\bar{z}_{n}-z^{*}),\Sigma^{1/2}Z\right)\coloneqq\sup_{h\in\mathrm{Lip}_{1}}\mathbb{E}\left[h\left(\sqrt{n}(\bar{z}_{n}-z^{*})\right)-h\left(\Sigma^{1/2}Z\right)\right].$ (14)
The asymptotic distribution of TSA-PR involves the Schur complement $G=A_{ff}-A_{fs}A_{ss}^{-1}A_{ss}$ of $A_{ss}$, as well as the Schur complement $\Delta$ defined in Section 2.
Theorem 3.2.
Let $\bar{z}_{n}\coloneqq(\bar{x}_{n},\bar{y}_{n})$ be the Ruppert-Polyak average of $(x_{t},y_{t})$ generated by (5). Define the asymptotic covariance of $\sqrt{n}(\bar{x}_{n}-x^{*})$ and $\sqrt{n}(\bar{y}_{n}-y^{*})$ as
$\bar{\Sigma}_{ff}=\lim_{n\to\infty}n\mathbb{E}(\bar{x}_{n}-x^{*})(\bar{x}_{n}-x^{*})^{T},\quad\bar{\Sigma}_{ss}=\lim_{n\to\infty}n\mathbb{E}(\bar{y}_{n}-y^{*})(\bar{y}_{n}-y^{*})^{T}.$ (15)
Under Assumptions 13, the above limits exist, and the rate of convergence in the Wasserstein-1 distance is given by:
$\begin{split}d_{1}\left(\sqrt{n}G(\bar{x}_{n}-x^{*}),(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(\frac{n^{a-1/2}}{a-1/2}+\frac{n^{(b-a)/2}}{b-a}+n^{b/2}\right)\right),\\ d_{1}\left(\sqrt{n}\Delta(\bar{y}_{n}-y^{*}),(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(\frac{n^{a-1/2}}{a-1/2}+\frac{n^{(b-a)/2}}{b-a}+n^{b/2}\right)\right),\end{split}$ (16)
where $Z_{1}$ and $Z_{2}$ are standard Gaussian vectors of appropriate dimensions.
A closed-form solution for $\bar{\Sigma}_{ff}$ and $\bar{\Sigma}_{ss}$ is given in Section C, Eq. (63). The error bound above suggests that an appropriate choice of step sizes is $a=1/2+1/\log n$ and $b=a+1/\log n$ for sufficiently large $n$. With this choice, the error bound is $\mathcal{O}(n^{-1/4})$.
By restricting the test functions $h$ in the Wasserstein-1 distance in Eq. (14) to be $h(x)\in\{-\lVert x\rVert,\lVert x\rVert\}$, we have the following lower and upper bounds on the expected error.
Corollary 3.3.
Let $Z_{1}$ and $Z_{2}$ be standard Gaussian vectors. Under Assumptions 13,
$\begin{split}\Big{\lvert}\mathbb{E}\lVert G(\bar{x}_{n}-x^{*})\rVert-\frac{1}{\sqrt{n}}\mathbb{E}\lVert(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\rVert\Big{\rvert}=o(n^{-1/2})\\ \Big{\lvert}\mathbb{E}\lVert\Delta(\bar{y}_{n}-y^{*})\rVert-\frac{1}{\sqrt{n}}\mathbb{E}\lVert(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\rVert\Big{\rvert}=o(n^{-1/2})\end{split}.$ (17)
The bounds on the expected error is tight, in the sense that the expected error is bounded above by rate $n^{-1/2}$, and similar to Eq. (13), both $\mathbb{E}\lVert\bar{x}_{n}-x^{*}\rVert$ and $\mathbb{E}\lVert\bar{y}_{n}-y^{*}\rVert$ are lower bounded by the expected norm of the limiting Gaussian vector times the rate $n^{-1/2}$. Replacing the norm of the Gaussian vector with the trace of its covariance would yield $\mathcal{O}(n^{-1/2})$ instead of the $o(n^{-1/2})$ above.
Remark 3.4.
As mentioned in Section 1.1, a finite-time performance guarantee on TSA-PR has not been established in literature. Second moment bounds, which are common metrics of interest in the analysis of stochastic approximation algorithms, are inevitably tied to the trace of an algorithm’s asymptotic covariance. We emphasize that such bounds only provide a conservative estimate of $n^{-1/2}\sqrt{\mathrm{Tr}\bar{\Sigma}}$ on the expected error.
3.1 Discussion
To understand the significance of the Wasserstein-1 distance error bounds achieved by TSA-PR, it is useful to compare with an oracle algorithm that has access to the system parameters $A$. Such an algorithm has been considered as a baseline in prior work (Polyak and Juditsky, 1992; Konda and Tsitsiklis, 2004; Mokkadem and Pelletier, 2006).
Consider a modification of TSA in (4), where the algorithm is allowed to design two gain matrices $Q_{ff},Q_{ss}$ and estimate the solution $(x^{*},y^{*})$ with the sequence $\{(x_{t}^{\dagger},y_{t}^{\dagger})\}$ updated as
$\begin{split}x_{t+1}^{\dagger}&=x_{t}^{\dagger}-\alpha_{t}Q_{ff}\left(A_{ff}x_{t}^{\dagger}+A_{fs}y_{t}^{\dagger}-W_{t}\right)\\ y_{t+1}^{\dagger}&=y_{t}^{\dagger}-\gamma_{t}Q_{ss}\left(A_{sf}x_{t}^{\dagger}+A_{ss}y_{t}^{\dagger}-V_{t}\right).\end{split}$ (18)
A constraint is imposed on $Q_{ff}$ and $Q_{ss}$ so that the modified system with the gain matrix is asymptotically stable. It was shown in (Mokkadem and Pelletier, 2006) that if the step sizes are chosen as $\alpha_{t}=\mathcal{O}(t^{-a})$ and $\gamma_{t}=\mathcal{O}(t^{-b})$ with $a\in(1/2,1)$, the slow-time-scale $\{y_{t}^{\dagger}\}_{t=1}^{n}$ converges in distribution to the solution $y^{*}$:
$\sqrt{n}(y_{n}^{\dagger}-y^{*})\to\mathcal{N}(0,\Sigma_{ss}(Q)),$
where $\Sigma_{ss}(Q_{ss})$ is the solution to the Lyapunov equation
$\begin{split}\left(Q_{ss}\Delta-\frac{1}{2}I\right)\Sigma_{ss}(Q_{ss})+\Sigma_{ss}(Q_{ss})\left(\Delta^{T}Q_{ss}^{T}-\frac{1}{2}I\right)\\ =Q_{ss}\left(\Gamma_{ss}+A_{sf}A_{ff}^{-1}\Gamma_{ff}(A_{sf}A_{ff}^{-1})^{T}-(A_{sf}A_{ff}^{-1}\Gamma_{fs}+(A_{sf}A_{ff}^{-1}\Gamma_{fs})^{T})\right)Q_{ss}^{T}.\end{split}$ (19)
A similar result holds for $x_{t}^{\dagger}$ when it is updated as the slow-time-scale variable (reversing the step size decay rates), where the other Schur complement $G$ appears in place of $\Delta$ in the above equation. The minimum mean square error and the corresponding asymptotic covariance matrix is obtained by solving
$\min_{Q_{ff}}\mathrm{Tr}\Sigma_{ff}(Q_{ff})\eqcolon\mathrm{Tr}\Sigma_{ff}^{*},\quad\min_{Q_{ss}}\mathrm{Tr}\Sigma_{ss}(Q_{ss})\eqcolon\mathrm{Tr}\Sigma_{ss}^{*},$
which can be shown to satisfy
$\begin{split}G\Sigma_{ff}^{*}G^{T}&=\Gamma_{ff}+A_{fs}A_{ss}^{-1}\Gamma_{ss}(A_{fs}A_{ss}^{-1})^{T}-\left(\Gamma_{fs}(A_{fs}A_{ss}^{-1})^{T}+A_{fs}A_{ss}^{-1}\Gamma_{sf}\right)\\ \Delta\Sigma_{ss}^{*}\Delta^{T}&=\Gamma_{ss}+A_{sf}A_{ff}^{-1}\Gamma_{ff}(A_{sf}A_{ff}^{-1})^{T}-\left(\Gamma_{sf}(A_{sf}A_{ff}^{-1})^{T}+A_{sf}A_{ff}^{-1}\Gamma_{fs}\right)\end{split}.$ (20)
This matches the asymptotic covariance achieved by TSA-PR. But in this case, the above covariances are not obtained simultaneously because one variable has to be updated at a faster time scale. This is in contrast to TSA-PR, which achieves the minimum asymptotic mean square errors for both time scales simultaneously and at the same rate. Next, we explain a gap in literature in the context of achievable rates.
Improved Rates via TSA-PR: Konda and Tsitsiklis (2004) proved that TSA converges asymptotically in m.s., where it was shown that
$\lim_{n\to\infty}\alpha_{n}^{-1}\mathbb{E}(x_{n}-x^{*})(x_{n}-x^{*})^{T}=\Sigma_{ff},\quad\lim_{n\to\infty}\gamma_{n}^{-1}\mathbb{E}(y_{n}-y^{*})(y_{n}-y^{*})^{T}=\Sigma_{ss},$ (21)
with $\Sigma_{ff},\Sigma_{ss}$ defined in Eq. (11). Finite-time counterparts to the above asymptotic convergence were established by Dalal et al. (2018); Kaledin et al. (2020); Haque et al. (2023), where it was shown that
$\mathbb{E}(x_{n}-x^{*})(x_{n}-x^{*})^{T}=\mathcal{O}\left(\alpha_{n}\right)\Sigma_{ff},\quad\mathbb{E}(y_{n}-y^{*})(y_{n}-y^{*})^{T}=\mathcal{O}\left(\gamma_{n}\right)\Sigma_{ss}.$ (22)
The MSE rates, obtained by taking the trace of the above, demonstrate that the asymptotic rates in Eq. (21) are achievable. Our upper bound on the norm of the difference in Theorem 3.1 is stronger than Eq. (22) in that it also proves a lower bound, namely that the mean square errors decay no faster than $\alpha_{n}$ and $\gamma_{n}$. These bounds suggest that the expected error rate achieved by TSA is of order $\mathcal{O}(n^{-a/2})$ and $\mathcal{O}(n^{-b/2})$, where $a,b$ satisfy Assumption 3.
The scaling factor of $\sqrt{n}$ in $\bar{x}_{n}-x^{*}$ and $\bar{y}_{n}-y^{*}$ in Theorem 3.2 suggests that TSA-PR enjoys faster rates of convergence than TSA. Moreover, its asymptotic covariance cannot be improved upon even when the model parameters are known, as seen for $(x_{n}^{\dagger},y_{n}^{\dagger})$. Therefore, we conclude that the $n^{-1/2}$ rate in Corollary 3.3 cannot be improved upon without a trade-off in covariance or other modes of convergences.
Next, we describe the tightness of the bound in Corollary 3.3. First, note that the result conveys both an upper and lower bound on $\mathbb{E}\lVert\bar{x}_{n}-x^{*}\rVert$ and $\mathbb{E}\lVert\bar{y}_{n}-y^{*}\rVert$, indicating that the guarantee is asymptotically tight. Knowing that the limit of $\sqrt{n}(\bar{x}_{n}-x^{*},\bar{y}_{n}-y^{*})$ is the Gaussian vector $\bar{\Sigma}^{1/2}Z$, the best possible bound on the second moment is of the form
$\mathbb{E}\lVert\bar{x}_{n}-x^{*}\rVert^{2}=\frac{1}{n}\mathrm{Tr}\bar{\Sigma}_{ff}+o(n^{-1}),\quad\mathbb{E}\lVert\bar{y}_{n}-y^{*}\rVert^{2}=\frac{1}{n}\mathrm{Tr}\bar{\Sigma}_{ss}+o(n^{-1}).$ (23)
Without additional information, the expected error can be deduced using Jensen’s inequality to obtain
$\mathbb{E}\lVert\bar{x}_{n}-x^{*}\rVert=\frac{1}{\sqrt{n}}\sqrt{\mathrm{Tr}\bar{\Sigma}_{ff}}+o(n^{-1/2}),\quad\mathbb{E}\lVert\bar{y}_{n}-y^{*}\rVert=\frac{1}{\sqrt{n}}\sqrt{\mathrm{Tr}\bar{\Sigma}_{ss}}+o(n^{-1/2}).$
But Corollary 3.3 establishes $\mathbb{E}\lVert\bar{x}_{n}-x^{*}\rVert=\Theta(n^{-1/2}\mathbb{E}\lVert\bar{\Sigma}_{ff}Z_{1}\rVert)$ and a similar improvement for $\mathbb{E}\lVert\bar{y}_{n}-y^{*}\rVert$. Therefore, it is interesting to see that the use of the Wassertein-1 distance in our CLT captures the expected error precisely up to exact constants.
4 Experiments
In this section, we complement our theoretical results with simulations that demonstrate the numerical observation of the finite-time bounds and CLT behaviors as predicted by our theories.
Figure 1: Errors achieved by TSA and TSA-PR when solving a synthetic system of equations with step sizes $\alpha_{n}=0.5/(n+1000)^{0.5}$, $\gamma_{n}=0.5/(n+1000)^{0.6}$. Left: fast-time-scale. Right: slow-time-scale. Errors are scaled by $\sqrt{n}$ to compare the rates, where the upward movement of TSA curves indicates that both fast- and slow-time-scales suffer a slower rate of convergence. TSA-PR achieves the $n^{-1/2}$ error rate.
Synthetic System of Equations. Our first set of experiments uses randomly generated system parameters $A_{ff},A_{fs},A_{sf},A_{ss}\in\mathbb{R}^{5\times 5}$ subject to the constraints in Assumption 1. The noise $\{(W_{t},V_{t})\}$ is sampled i.i.d. from the Gaussian distribution $\mathcal{N}(0,\Gamma)$ with $\Gamma_{fs}=0$ and $\Gamma_{ff},\Gamma_{ss}$ chosen to satisfy $\bar{\Sigma}_{ss}=I$ according to Eq. (20).
We implement TSA and TSA-PR with the step sizes $\alpha_{n}=0.5/(n+1000)^{0.5}$, $\gamma_{n}=0.5/(n+1000)^{0.6}$. By Corollary 3.3, the errors $\lVert\bar{x}_{n}-x^{*}\rVert,\lVert\bar{y}_{n}-y^{*}\rVert$ achieved by TSA-PR are expected to decay with rate $1/\sqrt{n}$, whereas Theorem 3.1 suggests that the errors $\lVert x_{n}-x^{*}\rVert,\lVert y_{n}-y^{*}\rVert$ achieved by TSA depend on the choice of step sizes and decay slower than $1/\sqrt{n}$. This is indeed observed in Figure 1, which plots the empirical errors scaled by $\sqrt{n}$, averaged over 100 trials. The TSA-PR curves for fast- and slow-time-scale variables exhibit downward movement, suggesting that $\|\bar{x}_{n}-x^{\star}\|$ and $\|\bar{y}_{n}-y^{\star}\|$ both decay with rate $\mathcal{O}(1/\sqrt{n})$. In comparison, the TSA curves move upward over iterations, which indicates that the iterates of TSA do not converge at an $1/\sqrt{n}$ rate.
According to Theorems 3.1 and 3.2, the scaled deviation $\sqrt{n}(x_{n}-x^{*},y_{n}-y^{*})$ of TSA exhibits increasing variance as $n$ grows, whereas the scaled deviation $\sqrt{n}(\bar{x}_{n}-x^{*},\bar{y}_{n}-y^{*})$ of TSA-PR converges to a Gaussian distribution. We visualize the deviations in Figure 2 using a smoothed kernel density plot at three checkpoints (different values of $n$). For simplicity of illustration, we only plot the first coordinates of the slow variables $\sqrt{n}(y_{n}-y^{*})$ and $\sqrt{n}(\bar{y}_{n}-y^{*}$). The figure reveals that $\sqrt{n}(\bar{y}_{n}-y^{*})$ indeed converges to a Gaussian distribution with standard deviation $1$, which is the asymptotic covariance predicted by Theorem 3.2. The distribution of $\sqrt{n}(y_{n}-y^{*})$ exhibits an increasing standard deviation over time, again matching the results in Theorem 3.1.
Figure 2: Gaussian behavior of slow-time-scale iterate around the solution $y^{*}$ with the choice of step sizes $\alpha_{n}=0.5/(n+1000)^{0.5}$, $\gamma_{n}=0.5/(n+1000)^{0.6}$. As $n$ increases, the variance $\sqrt{n}(y_{n,1}-y_{1}^{*})$ of TSA increases as suggested by Theorem 3.1, whereas that of TSA-PR converges and approaches 1.
Gradient-Based Policy Evaluation. Temporal difference learning with gradient correction (TDC) is a gradient-based policy evaluation algorithm in RL which works stably with function approximation and off-policy sampling (Sutton et al., 2009). The algorithm seeks to solve the system of equations
$\left\{\begin{array}[]{l}\mathbb{E}_{\pi}[\phi(s)\phi(s)^{\top}x-r(s,a)-\gamma\phi(s^{\prime})^{\top}y\phi(s)+\phi(s)^{\top}y\phi(s)]=0,\\ \mathbb{E}_{\pi}[\gamma\phi(s^{\prime})\phi(s)^{\top}x-r(s,a)-\gamma\phi(s^{\prime})^{\top}y\phi(s)+\phi(s)^{\top}y\phi(s)]=0.\end{array}\right.$ (26)
Here $x,y\in\mathbb{R}^{d}$ are the (fast- and slow-time-scale) variables, $s,a,s^{\prime}$ denote the state, action, and next state, and $\phi(s)\in\mathbb{R}^{d}$ is the feature vector associated with state $s$. Details on how TDC can be expressed in the form Eq. (4) can be found in (Xu et al., 2019; Zeng and Doan, 2024a).
We employ TSA (equivalent to the standard TDC method in this case) and TSA-PR to evaluate the cumulative reward of a random policy under linear function approximation and off-policy samples collected by a uniform behavior policy. The reward function, transition kernel, and feature vectors are all randomly generated, with $d=10$. Figure 3 compares the (scaled) errors of TSA and TSA-PR under the step sizes $\alpha_{n}=0.5/(n+1000)^{0.5}$, $\gamma_{n}=0.5/(n+1000)^{0.6}$, and again shows that the errors $\lVert\bar{x}_{n}-x^{*}\rVert,\lVert\bar{y}_{n}-y^{*}\rVert$ of TSA-PR converge with rate $1/\sqrt{n}$ while the errors $\lVert x_{n}-x^{*}\rVert,\lVert y_{n}-y^{*}\rVert$ of TSA do not converge as fast.
Figure 3: Scaled errors of TSA and TSA-PR for policy evaluation with the step sizes $\alpha_{n}=0.5/(n+1000)^{0.5}$, $\gamma_{n}=0.5/(n+1000)^{0.6}$. Left: fast-time-scale. Right: slow-time-scale. Again, the TSA curves move up over time, suggesting that both fast- and slow-time-scale errors decay at rate slower than $n^{-1/2}$. In comparison, TSA-PR achieves the $n^{-1/2}$ error rate.
5 Proof Sketches for Theorems 3.1 and 3.2
The analysis of two-time-scale algorithms is difficult because of the coupling between the fast- and slow-time-scale iterates. Methods for establishing finite-time MSE bounds on TSA are often based on a decoupling technique introduced by Konda and Tsitsiklis (2004). The complexity involved in the MSE analysis is further accentuated when analyzing the averaged iterates, because it involves the inner product between the averages which can be intractable. We observe that the Wasserstein-1 distance between errors achieved by TSA-PR and the limiting Gaussian variables can be made tractable by decomposing the TSA-PR errors as standardized martingale sums and weighted averages of the TSA following (Mokkadem and Pelletier, 2006). While Mokkadem and Pelletier (2006) prove asymptotic convergence by establishing that the TSA converges almost surely to zero and that the martingales converge in distribution, proving the rates of convergence requires a different approach. By considering the Wasserstein-1 distance as the metric, we exploit the useful property in Lemma 5.2 to bound the Wasserstein-1 distance with that of standardized martingale sums and terms that decay to zero in expected norm by Theorem 3.1. The martingale component converges in distribution to a non-zero limit, hence determines the limiting distribution. It is shown that the expected error achieved by TSA dominates the bound on the Wasserstein-1 distance.
Proof of Theorem 3.1: Our mean square analysis starts by using the decoupling technique from (Konda and Tsitsiklis, 2004), which is the cornerstone of two-time-scale analysis (Kaledin et al., 2020; Haque et al., 2023). A derivation of the details is provided in Section A.2. Recall the definition of the residues
$\hat{x}_{t}\coloneqq x_{t}-x_{\infty}(y_{t}),\quad\hat{y}_{t}\coloneqq y_{t}-y^{*},$ (27)
which measure the deviation from the steady-state solution $x_{\infty}(y)$ of the fast-time-scale system satisfying $F(x_{\infty}(y),y)=0$ for every $y$. Using a coordinate transformation
$\tilde{x}_{t}=\hat{x}_{t}+L_{t}\hat{y}_{t},$ (28)
the recursions for $(\tilde{x}_{t},\hat{y}_{t})$ can then be written as updates with respect to a new system $B$ obtained by a change of basis on the system $A$:
$\begin{split}\tilde{x}_{t+1}&=\tilde{x}_{t}-\alpha_{t}(B_{t}^{ff}\tilde{x}_{t}+B_{t}^{fs}\hat{y}_{t})+\alpha_{t}W_{t}+\gamma_{t}(L_{t+1}+A_{ff}^{-1}A_{fs})V_{t},\\ \hat{y}_{t+1}&=\hat{y}_{t}-\gamma_{t}(B_{t}^{sf}\tilde{x}_{t}+B_{t}^{ss}\hat{y}_{t})+\gamma_{t}V_{t},\end{split}$ (29)
where the specific expression for the transformed system $B$ is deferred to Sections B.1B.3. When $\{L_{t}\}$ is defined recursively with initial condition $L_{0}=0$ as
$L_{t+1}\coloneqq\left(L_{t}-\alpha_{t}A_{ff}L_{t}+\gamma_{t}A_{ff}^{-1}A_{fs}(\Delta-A_{sf}L_{t})\right)(I-\gamma_{t}(\Delta-A_{sf}L_{t}))^{-1},$ (30)
the sequence $\{B_{t}^{fs}\}$ in Eq. (29) is zero for every $t$, effectively removing the explicit dependency of the fast-time-scale $\tilde{x}_{t}$ from the slow-time-scale $\hat{y}_{t}$. A derivation for Eq. (30) is provided in Section A.2 The recursion is shown to be well defined in Proposition A.1, and the mean square error of the transformed iterate $\tilde{x}_{t}$ is then analyzed in Section B.1. Specifically, we show that for some constants $M_{f}^{\prime},M_{s}^{\prime}>0$,
$\begin{split}\lVert\mathbb{E}[\tilde{x}_{t+1}\tilde{x}_{t+1}^{T}]-\alpha_{t+1}\Sigma_{ff}\rVert&\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{2}+\gamma_{t}M_{f}^{\prime}\right)\lVert\mathbb{E}\tilde{x}_{t}\tilde{x}_{t}^{T}-\alpha_{t}\Sigma_{ff}\rVert+\mathcal{O}\left(\alpha_{t}\gamma_{t}\right),\\ \lVert\mathbb{E}[\hat{y}_{t+1}\hat{y}_{t+1}^{T}]-\gamma_{t+1}\Sigma_{ss}\rVert&\leq\left(1-\gamma_{t}\frac{\mu_{\Delta}}{2}+\frac{\gamma_{t}^{2}}{\alpha_{t}}M_{s}^{\prime}\right)\lVert\mathbb{E}\hat{y}_{t}\hat{y}_{t}^{T}-\gamma_{t}\Sigma_{ss}\rVert+\mathcal{O}\left(\gamma_{t}^{2}\right).\end{split}$ (31)
Once this is shown, a standard induction step can be applied with the choice of step sizes in Assumption 3 to obtain Theorem 3.1.
The second moment of $x_{t}-x^{*}$ is then recovered using the identities $x_{t}-x^{*}=\hat{x}_{t}+(x_{\infty}(y_{t})-x^{*})=\hat{x}_{t}+H\hat{y}_{t}$ to get
$\mathbb{E}\left(x_{t}-x^{*}\right)\left(x_{t}-x^{*}\right)^{T}=\mathbb{E}\tilde{x}_{t}\tilde{x}_{t}^{T}+(H-L_{t})\mathbb{E}\hat{y}_{t}\hat{y}_{t}^{T}(H-L_{t})^{T}+\mathbb{E}\tilde{x}_{t}\hat{y}_{t}^{T}(H-L_{t})^{T}.$
It is shown in Sections B.2 and B.3 that $\lVert\mathbb{E}\tilde{x}_{t}\hat{y}_{t}^{T}\rVert=\mathcal{O}(\gamma_{t})$ and $\lVert\mathbb{E}\hat{y}_{t}\hat{y}_{t}^{T}\rVert=\mathcal{O}(\gamma_{t})$. Proposition A.1 establishes that $\lVert L_{t}\rVert$ is uniformly bounded, from which we recover the mean square error of $x_{t+1}-x^{*}$ as
$\mathrm{Tr}\mathbb{E}(x_{t+1}-x^{*})(x_{t+1}-x^{*})\leq\alpha_{t}\mathrm{Tr}\Sigma_{ff}+\mathcal{O}\left(\gamma_{t}\right).$ (32)
This completes the mean square error bounds.
Proof of Theorem 3.2: Without loss of generality, let $x^{*}=0$ and $y^{*}=0$. In Section C, following Mokkadem and Pelletier (2006), we then express the PR averages as a telescoped sum
$\begin{split}G\bar{x}_{n}&=\frac{1}{n}\sum_{t=1}^{n}(W_{t}-A_{fs}A_{ss}^{-1}V_{t})+\frac{1}{n}\sum_{t=1}^{n}\alpha_{t}^{-1}(x_{t}-x_{t+1})-\frac{1}{n}\sum_{t=1}^{n}\gamma_{t}^{-1}A_{fs}A_{ss}^{-1}(y_{t}-y_{t+1})\\ \Delta\bar{y}_{n}&=\frac{1}{n}\sum_{t=1}^{n}(V_{t}-A_{sf}A_{ff}^{-1}W_{t})+\frac{1}{n}\sum_{t=1}^{n}A_{sf}A_{ff}^{-1}\alpha_{t}^{-1}(x_{t}-x_{t+1})+\frac{1}{n}\sum_{t=1}^{n}\gamma_{t}^{-1}(y_{t}-y_{t+1}).\end{split}$ (33)
After scaling Eq. (33) by $\sqrt{n}$, we see that the first terms are standardized sums of martingales which converge in distribution. To obtain their rates of convergence, we use the quantitative bound from (Srikant, 2024) on the Wasserstein-1 distance between a standardized sum of martingales and its limiting Gaussian variable.
Lemma 5.1.
Let $\{N_{t}\}$ be a martingale difference sequence in $\mathbb{R}^{d}$ satisfying $\mathbb{E}N_{t}N_{t}^{T}=\Gamma$ and $\mathbb{E}\lVert N_{t}\rVert^{2+\beta}<\infty$ for every $t\geq 1$ and some $\beta\in(0,1)$. For a standard Gaussian vector $Z$, it holds that
$d_{1}\left(n^{-1/2}\sum_{t=1}^{n}N_{t},\Gamma^{1/2}Z\right)\leq\mathcal{O}\left(\frac{d}{1-\beta}\right)\frac{\lVert\Gamma^{1/2}\rVert\lVert\Gamma^{-1/2}\rVert^{2+\beta}}{n^{\beta/2}}.$ (34)
Next, we relate the telescoped terms in Eq. (33) to their respective errors in expectation as
$\begin{split}\mathbb{E}\lVert\sum_{t=1}^{n}\alpha_{t}^{-1}(x_{t}-x_{t+1})\rVert&\leq\alpha_{1}^{-1}\mathbb{E}\lVert x_{1}\rVert+\alpha_{n}^{-1}\mathbb{E}\lVert x_{n}\rVert+\alpha_{1}^{-1}\sum_{t=1}^{n}t^{a-1}\mathbb{E}\lVert x_{t}\rVert\\ \mathbb{E}\lVert\sum_{t=1}^{n}\gamma_{t}^{-1}(y_{t}-y_{t+1})\rVert&\leq\gamma_{1}^{-1}\mathbb{E}\lVert y_{1}\rVert+\gamma_{n}^{-1}\mathbb{E}\lVert y_{n}\rVert+\gamma_{1}^{-1}\sum_{t=1}^{n}t^{b-1}\mathbb{E}\lVert y_{t}\rVert.\end{split}$ (35)
The expected errors of $x_{t}$ and $y_{t}$ for every $t\leq n$ are recovered from Theorem 3.1 and Jensen’s inequality; a detailed calculation is shown in Section C.
To combine these results, we then use the following property of the Wasserstein-1 distance.
Lemma 5.2.
Consider two random sequences $\{X_{t}\},\{Y_{t}\}$ such that $d_{1}(X_{t},X)\leq r_{t}$ for some random variable $X$ and $\mathbb{E}\lVert Y_{t}\rVert\leq r^{\prime}_{t}$. Then, $d_{1}(X_{t}+Y_{t},X)\leq r_{t}+r^{\prime}_{t}$.
Theorem 3.2 is then obtained by applying Lemma 5.2 to bound the distance between $G\bar{x}_{n}$ and its Gaussian limit with the Wasserstein-1 distance between the martingale terms in Eq. (33) and the expected norms in Eq. (35).
To obtain Corollary 3.3, we then use $h(x)=\lVert x\rVert$ as the 1-Lipschitz test function for the Wasserstein-1 distance in Eq. (14) to obtain
$\begin{split}\mathbb{E}\lVert G\bar{x}_{n}\rVert-\frac{1}{\sqrt{n}}\mathbb{E}\lVert(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\rVert&\leq\frac{1}{\sqrt{n}}d_{1}\left(\sqrt{n}G\bar{x}_{n},(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\right),\\ \mathbb{E}\lVert\Delta\bar{y}_{n}\rVert-\frac{1}{\sqrt{n}}\mathbb{E}\lVert(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\rVert&\leq\frac{1}{\sqrt{n}}d_{1}\left(\sqrt{n}\Delta\bar{y}_{n},(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\right).\end{split}$ (36)
From Theorem 3.2, the upper bound decays faster than $n^{1/2}$, and we have that $\mathbb{E}\lVert G\bar{x}_{n}\rVert$ and $\mathbb{E}\lVert\Delta\bar{y}_{n}\rVert$ are both $\mathcal{O}(n^{-1/2})$. The lower bound is obtained by using the test function $-\lVert x\rVert$.
6 Conclusion
We derived asymptotically tight finite-time bounds on the mean square error for the two-time-scale stochastic approximation, without PR averaging, driven by martingale noise. This was used to obtain finite-time bounds on the Wasserstein-1 distance between their Polyak-Ruppert averages and a Gaussian limit, which provides a non-asymptotic central limit theorem for TSA-PR. An important theoretical implication of our result is that it provides the first $n^{-1/2}$ finite-time bound on the expected error, which closes the gap between rates achievable for single- and two-time-scale stochastic approximation algorithms.
Acknowledgement
This work was supported by the National Science Foundation under awards ECCS-CAREER-2339509 and CCF-2343599.
Disclaimer
This paper was prepared for informational purposes [“in part” if the work is collaborative with external partners] by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates ("JP Morgan”) and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Appendix A Auxiliary Results
In this section, we first present the derivation for (29). This expression will be used to derive the finite-time bounds on the second moments of $\tilde{x}_{t}$ and $\hat{y}_{t}$. Some useful properties of recurring quantities will be derived, before presenting the proof of Theorems 3.1 and 3.2.
A.1 Recursion Updates
The goal of this sub-section is to derive the expressions for $B_{t}$ and $L_{t}$ in Eqs. (29) and (30), which is similar to the one in (Konda and Tsitsiklis, 2004). For convenience, recall the updates in (5)
$\begin{split}x_{t+1}&=x_{t}-\alpha_{t}\left(A_{ff}x_{t}+A_{fs}y_{t}-W_{t}\right),\\ y_{t+1}&=y_{t}-\gamma_{t}\left(A_{sf}x_{t}+A_{ss}y_{t}-V_{t}\right),\end{split}$
and the residues $\hat{x}_{t}\coloneqq x_{t}-x_{\infty}(y_{t})$ and $\hat{y}_{t}\coloneqq y_{t}-y^{*}$, where $x_{\infty}(y)=-A_{ff}^{-1}A_{fs}y$. The next steps will describe how to choose a sequence of matrices $\{L_{t}\}$ such that the recursion for $(\tilde{x}_{t+1},\hat{y}_{t+1})$ with $\tilde{x}_{t}\coloneqq\hat{x}_{t}+L_{t}\hat{y}_{t}$ can be simplified, and state the expressions for the resulting system.
Let $F,S$ be the operators corresponding to the fast- and slow-time-scale systems, i.e., $F(x,y)=A_{ff}x+A_{fs}y$ and $S(x,y)=A_{sf}x+A_{ss}y$. Using the definition of $x_{\infty}(y)$ we have
[…]
and we express $F$ and $S$ as a function of $\hat{x}_{t}$ and $\hat{y}_{t}$
$F(x_{t},y_{t})$ $=F(x_{t},y_{t})-F(x_{\infty}(y_{t}),y_{t})=A_{ff}\hat{x}_{t},$
$S(x_{t},y_{t})$ $=S(x_{t},y_{t})-S(x_{\infty}(y_{t}),y_{t})+S(x_{\infty}(y_{t}),y_{t})-S(x_{\infty}(y^{*}),y^{*})$
$=A_{sf}\hat{x}_{t}+A_{sf}x_{\infty}(\hat{y}_{t})+A_{ss}\hat{y}_{t}$
$=A_{sf}\hat{x}_{t}+\Delta\hat{y}_{t}.$
First, we obtain a recursion for the fast variable $\hat{x}_{t}$
$\hat{x}_{t+1}$ $=x_{t+1}-x_{\infty}(y_{t+1})$
$=\hat{x}_{t}-x_{\infty}(y_{t+1}-y_{t})-\alpha_{t}F(x_{t},y_{t})+\alpha_{t}W_{t}$
$=\hat{x}_{t}-\gamma_{t}A_{ff}^{-1}A_{fs}S(x_{t},y_{t})-\alpha_{t}A_{ff}\hat{x}_{t}+\alpha_{t}W_{t}-\gamma_{t}A_{ff}^{-1}A_{fs}V_{t}$
$=(I-\alpha_{t}A_{ff}-\gamma_{t}A_{ff}^{-1}A_{fs}A_{sf})\hat{x}_{t}-\gamma_{t}A_{ff}^{-1}A_{fs}\Delta\hat{y}_{t}+\alpha_{t}W_{t}-\gamma_{t}A_{ff}^{-1}A_{fs}V_{t}.$ (37)
Similarly, the slow-time-scale iterate satisfies the recursion
$\hat{y}_{t+1}$ $=y_{t+1}-y^{*}=\hat{y}_{t}-\gamma_{t}S(x_{t},y_{t})+\gamma_{t}V_{t}=(I-\gamma_{t}\Delta)\hat{y}_{t}-\gamma_{t}A_{sf}\hat{x}_{t}+\gamma_{t}V_{t}.$ (38)
Equations (37) and (38) can be combined to write the update rule for $\hat{z}_{t}=(\hat{x}_{t},\hat{y}_{t})$ in terms of a new system $\hat{A}_{t}$ and $Q_{t}$, where
$\hat{z}_{t+1}=\hat{z}_{t}-\hat{A}_{t}\hat{z}_{t}+Q_{t}N_{t},$
$\hat{A}_{t}=\begin{pmatrix}\alpha_{t}A_{ff}+\gamma_{t}A_{ff}^{-1}A_{fs}A_{sf}&\gamma_{t}A_{ff}^{-1}A_{fs}\Delta\\ \gamma_{t}A_{sf}&\gamma_{t}\Delta\end{pmatrix},\quad Q_{t}=\begin{pmatrix}\alpha_{t}I&-\gamma_{t}A_{ff}^{-1}A_{fs}\\ 0&\gamma_{t}I\end{pmatrix}.$
Second, we design a sequence of matrices $\{L_{t}\}$ such that the recursion for $\tilde{x}_{t}=\hat{x}_{t}+L_{t}\hat{y}_{t}$ does not depend on $\hat{y}_{t}$ explicitly. In the following, we use $P_{t}$ to denote a coordinate transformation matrix. The choice of $L_{t}$ determines the coordinate transformation
$\tilde{z}_{t}=P_{t}\hat{z}_{t},\quad P_{t}=\begin{pmatrix}I&L_{t}\\ 0&I\end{pmatrix}.$
For any $L_{t}$, it holds that
$\tilde{z}_{t+1}=P_{t+1}(I-\hat{A}_{t})P_{t}^{-1}\tilde{z}_{t}+P_{t+1}Q_{t}N_{t}.$ (39)
The matrix $L_{t}$ is to be chosen such that $P_{t+1}P_{t}^{-1}-P_{t+1}\hat{A}_{t}P_{t}^{-1}$ is lower triangular so that $\tilde{x}_{t+1}$ is decoupled from $\hat{y}_{t}$. Using $\hat{A}_{t}^{ff}$ and so on to denote the blocks in $\hat{A}_{t}$, we then have that the upper-right component of $P_{t+1}\hat{A}_{t}P_{t}^{-1}$ is given by
$L_{t+1}-L_{t}+(\hat{A}_{t}^{ff}+L_{t+1}\hat{A}_{t}^{sf})L_{t}-(\hat{A}_{t}^{fs}+L_{t+1}\hat{A}_{t}^{ss})=0.$
The above calculation is obtained by matrix multiplication and inverting $P_{t}$, which is equated to zero to achieve the decoupling. Substituting the expressions for $\hat{A}_{t}$ in the identity gives
$L_{t+1}(I+\hat{A}_{t}^{sf}L_{t}-\hat{A}_{t}^{ss})=\hat{A}_{t}^{fs}+(I-\hat{A}_{t}^{ff})L_{t},$
which yields the recursion for $L_{t}$ in (30). Given $L_{t}$, we finally derive the system matrix $B_{t}$ in Eq. (29). The expression for $B_{t}^{ff}$ can be easily obtained from Eq. (39), where $B_{t}^{fs}=0$ by design, as
$B_{t}^{ff}=A_{ff}+\frac{\gamma_{t}}{\alpha_{t}}(L_{t+1}+A_{ff}^{-1}A_{fs})A_{sf}.$
In addition, for any choice of $L_{t}$, $\hat{y}_{t+1}$ is related to $(\tilde{x}_{t},\hat{y}_{t})$ as follows:
$\hat{y}_{t+1}=(I-\gamma_{t}(\Delta-A_{sf}L_{t}))\hat{y}_{t}-\gamma_{t}A_{sf}\tilde{x}_{t}+\gamma_{t}V_{t},$
which yields the derivation for $B_{t}^{sf}=A_{sf}$ and $B_{t}^{ss}=\Delta-A_{sf}L_{t}$, where
$\hat{y}_{t+1}=\hat{y}_{t}-\gamma_{t}(B_{t}^{sf}\tilde{x}_{t}+B_{t}^{ss}\hat{y}_{t}-V_{t}).$ (40)
In summary, the choice of $L_{t}$ in Eq. (30) results in the updates
$\begin{split}\tilde{x}_{t+1}&=(I-\alpha_{t}A_{ff})\tilde{x}_{t}-\gamma_{t}\delta_{t}A_{sf}\tilde{x}_{t}+\alpha_{t}W_{t}+\gamma_{t}\delta_{t}V_{t},\\ \hat{y}_{t+1}&=(I-\gamma_{t}\Delta)\hat{y}_{t}+\gamma_{t}A_{sf}L_{t}\hat{y}_{t}-\gamma_{t}A_{sf}\tilde{x}_{t}+\gamma_{t}V_{t}.\end{split}$ (41)
A.2 Usage of Assumptions 13
Here we state a few elementary properties that will be used to prove Theorems 3.1 and 3.2. We will express the difference between finite-time second moments and the asymptotic covariance as a contraction. For a positive definite matrix $P$ and any matrix $A$, we define the weighted norm
$\lVert A\rVert^{2}_{P}=\sup_{x\neq 0}\frac{\langle Ax,Ax\rangle_{P}}{\langle x,x\rangle_{P}}=\sup_{x\neq 0}\frac{x^{T}A^{T}PAx}{x^{T}Px}.$
Note the equivalence of norms
$\lVert A\rVert\leq\sqrt{\kappa(P^{-1})}\lVert A\rVert_{P},\;\lVert A\rVert_{P}\leq\sqrt{\kappa(P)}\lVert A\rVert,$ (42)
where $\kappa(P)$ is the condition number of $P$. For a symmetric matrix $R$, observe
$\lVert R-\alpha(AR+RA^{T})\rVert_{P}$ $=\left\lVert\left(\frac{1}{2}I-\alpha A\right)R+R\left(\frac{1}{2}I-\alpha A^{T}\right)\right\rVert_{P}$
$\leq\lVert I-2\alpha A\rVert_{P}\lVert R\rVert_{P}.$
If $A$ satisfies Assumption 1, i.e., there exists a positive definite matrix $P$ such that
$AP+PA^{T}=I,$
then it is well known (e.g., Lemma 17 (Kaledin et al., 2020)) that for sufficiently small $\alpha$,
$\lVert I-2\alpha A\rVert_{P}\leq\sqrt{1-\frac{1}{2\lVert P\rVert^{2}}\alpha}\leq\left(1-\frac{\mu}{2}\alpha\right),$ (43)
where $\mu=(2\lVert P\rVert)^{-2}$, where the last inequality uses $\sqrt{1-x}\leq 1-x/2$. Hence, we conclude that the following contraction holds:
$\lVert R-\alpha(AR+RA^{T})\rVert_{P}\leq\left(1-\frac{\mu}{2}\right)\lVert R\rVert_{P}.$
Next, the choice of step size in Assumption 3 implies that
$\alpha_{t}-\alpha_{t+1}\leq\frac{\alpha_{t+1}}{t},$
obtained from the elementary inequality $(\frac{t+1}{t})^{a}\leq(1+\frac{1}{t})$ for every $a\in(0,1]$. Moreover, this implies $\alpha_{t+1}^{-1}-\alpha_{t}^{-1}\leq(\alpha_{t}t)^{-1}$ and is also true for the sequence $\{\gamma_{t}\}$, which is tight for the range of values $a,b$ we consider.
Finally, we derive a recursive bound for the absolute error of the second moments of fast and slow iterates from their asymptotic covariances. After an induction step, the dominant rates are then obtained using Lemma 14 (Kaledin et al., 2020), which states that for $\alpha_{1}\mu_{ff}/4\leq 1$ and constants
$\xi=1+\max\left\{\alpha_{1}\frac{\mu_{ff}}{4},\gamma_{1}\frac{\mu_{\Delta}}{16}\right\},\quad\xi^{\prime}=\frac{8}{\mu_{ff}}\max\left\{\xi,\frac{\mu_{ff}}{4\mu_{\Delta}}\xi,2\xi^{3}\right\},$
it holds that
$\begin{split}\sum_{t=1}^{n}\gamma_{t}\prod_{j=t+1}^{n}\left(1-\frac{\mu_{ff}}{4}\alpha_{j}\right)\leq\xi^{\prime}\frac{\gamma_{n}}{\alpha_{n}}\\ \sum_{t=1}^{n}\alpha_{t}\gamma_{t}\prod_{j=t+1}^{n}\left(1-\frac{\mu_{ff}}{4}\alpha_{j}\right)\leq\xi^{\prime}\gamma_{n},\\ \sum_{t=1}^{n}\frac{\alpha_{t}}{t}\prod_{j=t+1}^{n}\left(1-\frac{\mu_{ff}}{4}\alpha_{j}\right)\leq\xi^{\prime}\frac{1}{n}.\end{split}$ (44)
Variants of the above inequalities can also be deduced to be true, for example when $\gamma_{t}$ in the first inequality is replaced by $\gamma_{t}^{2}$.
A.3 Size of Linear Transformation
The following is from Lemma 18 and 19 in (Kaledin et al., 2020).
Proposition A.1.
The sequence $\{L_{t}\}$ is well-defined and converges to zero at rate $\mathcal{O}(\gamma_{t}/\alpha_{t})$.
The recursive definition for $L_{t+1}$ in Eq. (30) is equivalently expressed as
$L_{t+1}(I-\gamma_{t}\Delta+\gamma_{t}A_{sf}L_{t})$ $=\left(I-\alpha_{t}A_{ff}-\gamma_{t}A_{ff}^{-1}A_{fs}A_{sf}\right)L_{t}+\gamma_{t}A_{ff}^{-1}A_{fs}\Delta$
$=\left(I-\alpha_{t}A_{ff}\right)L_{t}+\gamma_{t}A_{ff}^{-1}A_{fs}(\Delta-A_{sf}L_{t}).$
Denoting $B_{t}^{ss}=\Delta-A_{sf}L_{t}$ as before,
$L_{t+1}(I-\gamma_{t}B_{t}^{ss})=(I-\alpha_{t}A_{ff})L_{t}+\gamma_{t}A_{ff}^{-1}A_{fs}B_{t}^{ss}.$ (45)
A contraction occurs for $L_{t+1}$ in the norm
$\lVert L\rVert_{P_{\Delta},P_{ff}}^{2}=\sup_{x\neq 0}\frac{\langle Lx,Lx\rangle_{P_{ff}}}{\langle x,x\rangle_{P_{\Delta}}},$
where it was shown that
$\lVert L_{t+1}\rVert_{P_{\Delta},P_{ff}}\leq\left(1-\frac{\mu_{ff}}{2}\alpha_{t}\right)\lVert L_{t}\rVert_{P_{\Delta},P_{ff}}+\mathcal{O}(\gamma_{t}).$
By induction and Eq. (44),
$\lVert L_{t+1}\rVert_{P_{\Delta},P_{ff}}$ $\leq\prod_{t=1}^{n}\left(1-\frac{\mu_{ff}}{2}\alpha_{t}\right)\lVert L_{1}\rVert_{P_{\Delta},P_{ff}}+\mathcal{O}\left(\frac{\gamma_{n}}{\alpha_{n}}\right).$
The decay in the unweighted norm holds at the same rate by equivalence of norms.
Appendix B Mean Square Analysis of TSA
In this section, we use the above properties to prove Theorem 3.1. The proof is divided into three parts, starting from the second moment of $\tilde{x}_{t}$, the joint covariance $\mathbb{E}\tilde{x}_{t}\hat{y}_{t}$, and the second moment of the slow-time-scale $\hat{y}_{t}$. Only the second moments of $\tilde{x}_{t}$ and $\hat{y}_{t}$ are included in the statement of Theorem 3.1, but the finite-time bound on the joint covariance is required as an intermediate step to establish the second moment of the slow-time-scale.
The transformed fast-time-scale $\tilde{x}_{t}$ is first analyzed because it is expressed recursively without an explicit dependence on $\hat{y}_{t}$. This will be used to derive the finite-time bounds on the joint-time-scale covariance $\mathbb{E}\tilde{x}_{t}\hat{y}_{t}^{T}$, which is used as an intermediate result to derive the finite-time bound on the slow-time-scale covariance $\mathbb{E}\hat{y}_{t}\hat{y}_{t}^{T}$.
B.1 Fast-Time-Scale Mean-Square Convergence
Consider the transformed last iterate $\tilde{x}_{t}=\hat{x}_{t}+L_{t}\hat{y}_{t}$
$\tilde{x}_{t+1}=\tilde{x}_{t}-\alpha_{t}B_{t}^{ff}\tilde{x}_{t}+\alpha_{t}W_{t}+\gamma_{t}(L_{t+1}+A_{ff}^{-1}A_{fs})V_{t},$ (46)
where $\hat{x}_{t}=x_{t}-x_{\infty}(y_{t}),\hat{y}_{t}=y_{t}-y^{*}$, and
$B_{t}^{ff}=A_{ff}+\frac{\gamma_{t}}{\alpha_{t}}(L_{t+1}+A_{ff}^{-1}A_{fs})A_{sf}.$ (47)
The quantity $L_{t+1}+A_{ff}^{-1}A_{fs}$ appears frequently, so we refer to it as $\delta_{t}$ stated here for future reference:
$\delta_{t}=L_{t+1}+A_{ff}^{-1}A_{fs}.$ (48)
We use $X_{t}$ to denote $\mathbb{E}[\tilde{x}_{t}\tilde{x}_{t}^{T}]$. In addition, let $\Sigma_{ff}$ be the unique solution to
$A_{ff}\Sigma_{ff}+\Sigma_{ff}A_{ff}^{T}=\Gamma_{ff}.$ (49)
We note that $\Sigma_{ff}$ exists since $-A_{ff}$ is Hürwitz.
Lemma B.1.
Under Assumptions 1-3, there exists a problem-dependent constant $M_{f}>0$ such that
$\lVert X_{n+1}-\alpha_{n+1}\Sigma_{ff}\rVert\leq\kappa(P_{ff})\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert\mathbb{E}X_{1}-\alpha_{1}\Sigma_{ff}\rVert+M_{f}\gamma_{n}.$
In particular,
$\mathbb{E}\lVert\tilde{x}_{n}\rVert^{2}=\alpha_{n}\mathrm{Tr}\Sigma_{ff}+\mathcal{O}\left(\gamma_{n}\right).$ (50)
Proof: We start by evaluating the expectation of $\tilde{x}_{t}\tilde{x}_{t}^{T}$ from (46) conditioned on $\mathcal{H}_{t}$:
$\mathbb{E}[\tilde{x}_{t+1}\tilde{x}_{t+1}|\mathcal{H}_{t}]$ $=\tilde{x}_{t}\tilde{x}_{t}^{T}-\alpha_{t}\left(B_{t}^{ff}\tilde{x}_{t}\tilde{x}_{t}^{T}+\tilde{x}_{t}\tilde{x}_{t}^{T}(B_{t}^{ff})^{T}-\alpha_{t}\Gamma_{ff}\right)$
$+\gamma_{t}^{2}\delta_{t}\Gamma_{ss}\delta_{t}^{T}+\alpha_{t}\gamma_{t}(\Gamma_{fs}\delta_{t}^{T}+\delta_{t}\Gamma_{sf}).$
Taking the unconditional expectations, we then obtain
$X_{t+1}=X_{t}-\alpha_{t}(B_{t}^{ff}X_{t}+X_{t}(B_{t}^{ff})^{T}-\alpha_{t}\Gamma_{ff})+\gamma_{t}^{2}\delta_{t}\Gamma_{ss}\delta_{t}^{T}+\alpha_{t}\gamma_{t}(\Gamma_{fs}\delta_{t}^{T}+\delta_{t}\Gamma_{sf})$
Using the expression for $B_{t}^{ff}$ in Eq. (47) into the preceeding equation gives
$X_{t+1}$ $=X_{t}-\alpha_{t}(A_{ff}X_{t}^{T}+X_{t}A_{ff}^{T}-\alpha_{t}\Gamma_{ff}),+\gamma_{t}R_{t}^{ff}(X_{t})$
$R_{t}^{ff}(X_{t})$ $\coloneqq-\left(\delta_{t}A_{sf}X_{t}+X_{t}A_{sf}^{T}\delta_{t}^{T}\right)+\left(\gamma_{t}\delta_{t}\Gamma_{ss}\delta_{t}^{T}+\alpha_{t}\Gamma_{fs}\delta_{t}^{T}+\alpha_{t}\delta_{t}\Gamma_{sf}\right),$
which by using the expression in Eq. (49) for $\Gamma_{ff}$ yields
$X_{t+1}-\alpha_{t+1}\Sigma_{ff}$ $=X_{t}-\alpha_{t}\Sigma_{ff}-\alpha_{t}\left(A_{ff}(X_{t}-\alpha_{t}\Sigma_{ff})+(X_{t}-\alpha_{t}\Sigma_{ff})A_{ff}^{T}\right)$
$+(\alpha_{t}-\alpha_{t+1})\Sigma_{ff}+\gamma_{t}R_{t}^{ff}(X_{t}).$ (51)
Define $M_{f}^{\prime}$ to be a constant such that
$\lVert R_{t}^{ff}(X_{t})\rVert_{P_{ff}}\leq M_{f}^{\prime}(\lVert X_{t}\rVert_{P_{ff}}+\alpha_{t})\leq M_{f}^{\prime}\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert_{P_{ff}}+M_{f}^{\prime}\alpha_{t}(1+\lVert\Sigma_{ff}\rVert_{P_{ff}}).$
Such a constant exists because $\lVert\delta_{t}\rVert$ in (48) is uniformly bounded by Proposition A.1. The choice of step sizes ensures Eq. (43) is satisfied for the symmetric matrix $X_{t}-\alpha_{t}\Sigma_{ff}$. By the property $\alpha_{t}-\alpha_{t+1}\leq t^{-1}\alpha_{t+1}$,
$\lVert X_{t+1}-\alpha_{t+1}\Sigma_{ff}\rVert_{P_{ff}}$ $\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{2}+\gamma_{t}M_{f}^{\prime}\right)\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert_{P_{ff}}+\frac{\alpha_{t+1}}{t}\lVert\Sigma_{ff}\rVert_{P_{ff}}$
$+\alpha_{t}\gamma_{t}M_{f}^{\prime}(\lVert\Sigma_{ff}\rVert_{P_{ff}}+1)$
$\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert_{P_{ff}}+\frac{\alpha_{t}}{t}\lVert\Sigma_{ff}\rVert_{P_{ff}}$
$+\alpha_{t}\gamma_{t}M_{f}^{\prime}(\lVert\Sigma_{ff}\rVert_{P_{ff}}+1).$
The preceding equation is of the form
$\lVert X_{t+1}-\alpha_{t+1}\Sigma_{ff}\rVert_{P_{ff}}\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert_{P_{ff}}+\mathcal{O}\left(\alpha_{t}\gamma_{t}+\frac{\alpha_{t}}{t}\right).$ (52)
By induction and Eq. (44), we have that for some constant $M_{f}>0$, it holds that
$\lVert X_{n+1}-\alpha_{n+1}\Sigma_{ff}\rVert_{P_{ff}}$ $\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{1}-\alpha_{1}\Sigma_{ff}\rVert_{P_{ff}}$
$+M_{f}^{\prime\prime}\sum_{t=1}^{n}\alpha_{t}\left(\gamma_{t}+\frac{1}{t}\right)\prod_{j=t+1}^{n}\left(1-\alpha_{j}\frac{\mu_{ff}}{4}\right)$
$\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{1}-\alpha_{1}\Sigma_{ff}\rVert_{P_{ff}}+M_{f}\left(\gamma_{n}+\frac{1}{n}\right)$
Using equivalence of norms, we obtain a bound on the unweighted norm
$\lVert X_{n+1}-\alpha_{n+1}\Sigma_{ff}\rVert$ $\leq\sqrt{\kappa(P_{ff}^{-1})}\lVert X_{n+1}-\alpha_{n+1}\Sigma_{ff}\rVert_{P_{ff}}$
$\leq\kappa(P_{ff})\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{1}-\alpha_{1}\Sigma_{ff}\rVert+\mathcal{O}\left(\gamma_{n}+\frac{1}{n}\right),$
where we re-use the constant $M_{f}$ in the result. By substituting $\hat{x}_{t}=\tilde{x}_{t}-L_{t}\hat{y}_{t}$, we obtain a bound for $\lVert\mathbb{E}\hat{x}_{t}\hat{x}_{t}^{T}-\alpha_{t+1}\Sigma_{ff}\rVert$. A detailed calculation can be found in Section C.
B.2 Joint-Time-Scale Mean Square Convergence
The analysis of the slow iterate’s covariance $\mathbb{E}\hat{y}_{t}\hat{y}_{t}^{T}$ involves an error term that depends on the deviation of cross covariance $C_{t}=\mathbb{E}[\tilde{x}_{t}\hat{y}_{t}^{T}]$ from $\gamma_{t}\Sigma_{fs}$. Because $x,y$ can have different dimensions, the contraction is proved for the cross covariance in the norm
$\lVert C\rVert_{P_{\Delta},P_{ff}}^{2}=\sup_{v\neq 0}\frac{v^{T}C^{T}P_{ff}Cv}{v^{T}P_{\Delta}v}.$
Recall the recursions in Eq. (41) derived in Section A.1, restated for reference below:
$\tilde{x}_{t+1}$ $=(I-\alpha_{t}A_{ff})\tilde{x}_{t}-\gamma_{t}\delta_{t}A_{sf}\tilde{x}_{t}+\alpha_{t}W_{t}+\gamma_{t}\delta_{t}V_{t}$
$\hat{y}_{t+1}$ $=(I-\gamma_{t}\Delta)\hat{y}_{t}+\gamma_{t}A_{sf}L_{t}\hat{y}_{t}-\gamma_{t}A_{sf}\tilde{x}_{t}+\gamma_{t}V_{t}.$
This section is dedicated to proving the following intermediate result in Theorem 3.1:
Lemma B.2.
Let $\Sigma_{fs}$ be the unique solution to
$A_{ff}\Sigma_{fs}+\Sigma_{ff}A_{sf}^{T}=\Gamma_{fs}.$ (53)
Under Assumptions 13, there is a problem-dependent constant $M_{fs}>0$ such that
$\lVert C_{n+1}-\gamma_{n+1}\Sigma_{fs}\rVert\leq\sqrt{\kappa(P_{ff})\kappa(P_{\Delta})}\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert+M_{fs}\frac{\gamma_{n}^{2}}{\alpha_{n}}$ (54)
Proof: Using the expressions in (41), we obtain that $C_{t+1}=\mathbb{E}\tilde{x}_{t+1}\hat{y}_{t+1}^{T}$ is given by
$C_{t+1}$ $=(I-{\color[rgb]{0,0,0}\alpha_{t}A_{ff}})C_{t}-\gamma_{t}C_{t}\Delta^{T}+\alpha_{t}\gamma_{t}A_{ff}C_{t}\Delta^{T}+\gamma_{t}\left(C_{t}L_{t}^{T}A_{sf}^{T}-{\color[rgb]{0,0,0}X_{t}A_{sf}^{T}}\right)$
$-\alpha_{t}\gamma_{t}A_{ff}\left(C_{t}L_{t}^{T}A_{sf}^{T}-X_{t}A_{sf}^{T}\right)-\gamma_{t}\delta_{t}A_{sf}C_{t}+\gamma_{t}^{2}\delta_{t}A_{sf}\left(C_{t}\Delta^{T}+C_{t}L_{t}^{T}A_{sf}^{T}-X_{t}A_{sf}^{T}\right)$
$-\alpha_{t}{\color[rgb]{0,0,0}\gamma_{t}\Gamma_{fs}}+\gamma_{t}^{2}\delta_{t}\Gamma_{ss}.$
Grouping terms by their rates (i.e., order-dependency on $t$ as determined by the step sizes),
$C_{t+1}$ $=C_{t}-\gamma_{t}\left(\frac{\alpha_{t}}{\gamma_{t}}A_{ff}C_{t}+X_{t}A_{sf}^{T}-\alpha_{t}\Gamma_{fs}\right)-\gamma_{t}\delta_{t}\left(A_{sf}C_{t}+C_{t}\Delta^{T}-\gamma_{t}\Gamma_{ss}\right)$
$+\alpha_{t}\gamma_{t}A_{ff}X_{t}A_{sf}^{T}+\gamma_{t}\left(\alpha_{t}A_{ff}C_{t}\Delta^{T}-\gamma_{t}\delta_{t}A_{sf}X_{t}A_{sf}^{T}\right)$
$-\gamma_{t}\left(\alpha_{t}A_{ff}C_{t}L_{t}^{T}A_{sf}^{T}-\gamma_{t}\delta_{t}A_{sf}C_{t}\Delta^{T}\right)+\gamma_{t}C_{t}L_{t}^{T}A_{sf}^{T}+\gamma_{t}^{2}\delta_{t}A_{sf}C_{t}L_{t}^{T}A_{sf}^{T}$
$=C_{t}-\gamma_{t}\left(\frac{\alpha_{t}}{\gamma_{t}}A_{ff}C_{t}+X_{t}A_{sf}^{T}-\alpha_{t}\Gamma_{fs}\right)+\gamma_{t}R_{t}^{fs}(C_{t}),$
where $R_{t}^{fs}(C_{t})$ subsumes all remaining terms. Substituting for $\Gamma_{fs}$ using Eq. (11),
$C_{t+1}$ $=\left(I-\alpha_{t}A_{ff}\right)C_{t}-\gamma_{t}\left(X_{t}A_{sf}^{T}-\alpha_{t}\Gamma_{fs}\right)+\gamma_{t}R_{t}^{fs}(C_{t})$
$=(I-\alpha_{t}A_{ff})C_{t}-\gamma_{t}\left((X_{t}-\alpha_{t}\Sigma_{ff})A_{sf}^{T}-\alpha_{t}A_{ff}\Sigma_{fs}\right)+\gamma_{t}R_{t}^{fs}(C_{t})$
$=C_{t}-\alpha_{t}A_{ff}(C_{t}-\gamma_{t}\Sigma_{fs})-\gamma_{t}\left(X_{t}-\alpha_{t}\Sigma_{ff}\right)A_{sf}^{T}+\gamma_{t}R_{t}^{fs}(C_{t}).$ (55)
Subtracting $\gamma_{t+1}\Sigma_{fs}$ on both sides,
$C_{t+1}-\gamma_{t+1}\Sigma_{fs}$
$=(I-\alpha_{t}A_{ff})(C_{t}-\gamma_{t}\Sigma_{fs})+(\gamma_{t}-\gamma_{t+1})\Sigma_{fs}-\gamma_{t}(X_{t}-\alpha_{t}\Sigma_{ff})A_{sf}^{T}+\gamma_{t}R_{t}^{fs}(C_{t}).$
Using Lemma B.1 to get $\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert=\mathcal{O}(\gamma_{t})$, and defining a constant $M_{fs}^{\prime}$ such that
$\lVert R_{t}^{fs}(C_{t})\rVert_{P_{\Delta},P_{ff}}\leq M_{fs}^{\prime}\left(\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+\alpha_{t}\lVert X_{t}\rVert+\gamma_{t}\lVert\Sigma_{fs}\rVert\right)$
which exists by sub-multiplicativity of operator norms, we then have for $\alpha_{t}\mu_{ff}/2-\gamma_{t}M_{fs}^{\prime}\geq\mu_{ff}/4$
$\lVert C_{t+1}-\gamma_{t+1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}$ $\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{2}+\gamma_{t}M_{fs}^{\prime}\right)\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+\frac{\gamma_{t}}{t}\lVert\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}$
$+\gamma_{t}\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert_{P_{ff}}\lVert A_{sf}\rVert_{P_{\Delta},P_{ff}}+\gamma_{t}M_{fs}\left(\alpha_{t}\lVert X_{t}\rVert+\gamma_{t}\lVert\Sigma_{fs}\rVert\right).$
$\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+M_{fs}^{\prime\prime}\left(\frac{\gamma_{t}}{t}\lVert\Sigma_{fs}\rVert+\gamma_{t}^{2}\lVert A_{sf}\rVert+\alpha_{t}^{2}\gamma_{t}\lVert\Sigma_{ff}\rVert\right)$
for some constant $M_{fs}^{\prime\prime}$. Since $b<1$ and $\alpha_{t}^{2}\leq\gamma_{t}$, $\gamma_{t}^{2}$ dominates and the above is a recursion of the form
$\lVert C_{t+1}-\gamma_{t+1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+M_{fs}^{\prime}\gamma_{t}^{2}$ (56)
for some constant $M_{fs}^{\prime}>0$. By induction and Eq. (44), we have that for a constant $M_{fs}>0$,
$\lVert C_{n+1}-\gamma_{n+1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}$ $\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+M_{fs}^{\prime}\sum_{t=1}^{n}\gamma_{t}^{2}\prod_{j=t+1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)$
$\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+M_{fs}\frac{\gamma_{n}^{2}}{\alpha_{n}}.$
The unweighted norm is recovered by the equivalence of norms
$\lVert C_{n+1}-\gamma_{n+1}\Sigma_{fs}\rVert$ $\leq\sqrt{\frac{\lambda_{\max}(P_{ff}^{-1})}{\lambda_{\min}(P_{\Delta}^{-1})}}\lVert C_{n+1}-\gamma_{n+1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}$
$\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}$ $\leq\sqrt{\frac{\lambda_{\max}(P_{ff})}{\lambda_{\min}(P_{\Delta})}}\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert.$
B.3 Slow-Time-Scale Mean Square Convergence
The bounds above for fast- and cross-covariance matrices $X_{n}$ and $C_{n}$ are used to derive the convergence rate of $Y_{n}=\mathbb{E}[\hat{y}_{n}\hat{y}_{n}^{T}]$.
Lemma B.3.
Let $\Sigma_{ss}\succ 0$ be the unique solution to the equation
$\Delta\Sigma_{ss}+\Sigma_{ss}\Delta^{T}=\Sigma_{fs}A_{fs}^{T}+A_{fs}\Sigma_{fs}^{T}+\Gamma_{ss}.$ (57)
Under Assumptions 13, there is a problem-dependent constant $M_{s}>0$ such that
$\lVert Y_{n+1}-\gamma_{n+1}\Sigma_{ss}\rVert\leq\kappa(P_{\Delta})\prod_{t=1}^{n}\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert Y_{1}-\gamma_{1}\Sigma_{ss}\rVert+\frac{M_{s}}{n}.$
Start from the same slow iterate’s recursion as above:
$\hat{y}_{t+1}=(I-\gamma_{t}\Delta)\hat{y}_{t}+\gamma_{t}A_{sf}L_{t}\hat{y}_{t}-\gamma_{t}A_{sf}\tilde{x}_{t}+\gamma_{t}V_{t}.$ (58)
The recursion for $Y_{t+1}$ can be derived by the tower property of expectations, where we take the conditional expectation with respect to $\mathcal{H}_{t}$ and then the unconditional expectation as done for $X_{t+1}$ in Section B.1:
$Y_{t+1}$ $=Y_{t}-\gamma_{t}{\color[rgb]{0,0,1}(\Delta Y_{t}+Y_{t}\Delta^{T})}+\gamma_{t}^{2}(\Delta Y_{t}\Delta^{T})$
$+\gamma_{t}\left((I-\gamma_{t}\Delta)Y_{t}L_{t}^{T}A_{sf}^{T}+((I-\gamma_{t}\Delta)Y_{t}L_{t}^{T}A_{sf}^{T})^{T}\right)$
$-\gamma_{t}\left(({\color[rgb]{0,0,1}I}-\gamma_{t}\Delta)C_{t}^{T}A_{sf}^{T}+(({\color[rgb]{0,0,1}I}-\gamma_{t}\Delta)C_{t}^{T}A_{sf}^{T})^{T}\right)$
$+\gamma_{t}^{2}A_{sf}L_{t}Y_{t}L_{t}^{T}A_{sf}^{T}-\gamma_{t}^{2}\left(A_{sf}L_{t}C_{t}^{T}A_{sf}^{T}+(A_{sf}L_{t}C_{t}^{T}A_{sf}^{T})^{T}\right)$
$+\gamma_{t}^{2}A_{sf}X_{t}A_{sf}^{T}+{\color[rgb]{0,0,1}\gamma_{t}^{2}\Gamma_{ss}}.$
Grouping terms by magnitude as done previously,
$Y_{t+1}$ $=Y_{t}-\gamma_{t}\left(\Delta Y_{t}+Y_{t}\Delta^{T}-(C_{t}^{T}A_{sf}^{T}+A_{sf}C_{t})-\gamma_{t}\Gamma_{ss}\right)+\frac{\gamma_{t}^{2}}{\alpha_{t}}R_{t}^{ss}(Y_{t}),$
where the remainder terms $R_{t}^{ss}$ is affine in $Y_{t}$. Let $M_{s}^{\prime}>0$ be a constant such that
$\lVert R_{t}^{ss}(Y_{t})\rVert_{P_{\Delta}}\leq M_{s}^{\prime}\left(\lVert Y_{t}\rVert_{P_{\Delta}}+\alpha_{t}\lVert A_{sf}C_{t}\rVert_{P_{\Delta}}+\alpha_{t}\lVert X_{t}\rVert_{P_{\Delta}}\right).$
We use $\lVert C_{t}\rVert=\mathcal{O}(\gamma_{t})$ and $\lVert X_{t}\rVert=\mathcal{O}(\alpha_{t})$ from Lemmas B.1 and B.2 to deduce a constant $M_{s}^{\prime\prime}$ such that
$\lVert R_{t}^{ss}(Y_{t})\rVert_{P_{\Delta}}\leq M_{s}^{\prime\prime}\left(\lVert Y_{t}\rVert_{P_{\Delta}}+\alpha_{t}^{2}\right)\leq M_{s}^{\prime\prime}\left(\lVert Y_{t}-\gamma_{t}\Sigma_{ss}\rVert_{P_{\Delta}}+\gamma_{t}\lVert\Sigma_{ss}\rVert_{P_{\Delta}}+\alpha_{t}^{2}\right)$
Using the definition of $\Sigma_{ss}$ that solves Eq. (57) to substitute for $\Gamma_{ss}$,
$Y_{t+1}$ $=Y_{t}-\gamma_{t}\left(\Delta Y_{t}+Y_{t}\Delta^{T}-\gamma_{t}(\Sigma_{fs}^{T}A_{sf}^{T}+A_{sf}\Sigma_{fs})-\gamma_{t}\Gamma_{ss}\right)$
$+\gamma_{t}\left((C_{t}-\gamma_{t}\Sigma_{fs})^{T}A_{sf}^{T}+A_{sf}(C_{t}-\gamma_{t}\Sigma_{fs})\right)+\frac{\gamma_{t}^{2}}{\alpha_{t}}R_{t}^{ss}(Y_{t})$
$=Y_{t}-\gamma_{t}\left(\Delta(Y_{t}-\gamma_{t}\Sigma_{ss})+(Y_{t}-\gamma_{t}\Sigma_{ss})\Delta^{T}\right)$
$+\gamma_{t}\left((C_{t}-\gamma_{t}\Sigma_{fs})^{T}A_{sf}^{T}+A_{sf}(C_{t}-\gamma_{t}\Sigma_{fs})\right)+\frac{\gamma_{t}^{2}}{\alpha_{t}}R_{t}^{ss}(Y_{t}).$
Subtracting both sides by $\gamma_{t+1}\Sigma_{ss}$,
$Y_{t+1}-\gamma_{t+1}\Sigma_{ss}$ $=Y_{t}-\gamma_{t}\Sigma_{ss}-\gamma_{t}\left(\Delta(Y_{t}-\gamma_{t}\Sigma_{ss})+(Y_{t}-\gamma_{t}\Sigma_{ss})\Delta^{T}\right)+(\gamma_{t}-\gamma_{t+1})\Sigma_{ss}$
$+\gamma_{t}\left((C_{t}-\gamma_{t}\Sigma_{fs})^{T}A_{sf}^{T}+A_{sf}(C_{t}-\gamma_{t}\Sigma_{fs})\right)+\frac{\gamma_{t}^{2}}{\alpha_{t}}R_{t}^{ss}(Y_{t}).$ (59)
Using Eq. (43) and $\gamma_{t}-\gamma_{t+1}\leq\gamma_{t}/t$,
$\lVert Y_{t+1}-\gamma_{t+1}\Sigma_{ss}\rVert_{P_{\Delta}}$ $\leq\left(1-\gamma_{t}\frac{\mu_{\Delta}}{2}+\frac{\gamma_{t}^{2}}{\alpha_{t}}M_{s}^{\prime}\right)\lVert Y_{t}-\gamma_{t}\Sigma_{ss}\rVert_{P_{\Delta}}+\frac{\gamma_{t}}{t}\lVert\Sigma_{ss}\rVert_{P_{\Delta}}$
$+\gamma_{t}\left\lVert(C_{t}-\gamma_{t}\Sigma_{fs})^{T}A_{sf}^{T}+A_{sf}(C_{t}-\gamma_{t}\Sigma_{fs})\right\rVert_{P_{\Delta}}+\frac{\gamma_{t}^{2}}{\alpha_{t}}M_{s}^{\prime}\left(\gamma_{t}\lVert\Sigma_{ss}\rVert_{P_{\Delta}}+\alpha_{t}^{2}\right).$
The choice of step size (10) ensures a contraction, and Lemma B.2 implies $\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert=\mathcal{O}(\gamma_{t}^{2}/\alpha_{t})$. Using that the above equation is of the form
$\lVert Y_{t+1}-\gamma_{t+1}\Sigma_{ss}\rVert_{P_{\Delta}}\leq\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert Y_{t}-\gamma_{t}\Sigma_{ss}\rVert_{P_{\Delta}}+\frac{\gamma_{t}}{t}M_{s}^{\prime\prime},$ (60)
we have by induction and Eq. (44) that for some constant $M_{s}$,
$\lVert Y_{n+1}-\gamma_{n+1}\Sigma_{ss}\rVert\leq\kappa(P_{\Delta})\prod_{t=1}^{n}\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert Y_{1}-\gamma_{1}\Sigma_{ss}\rVert+\frac{M_{s}}{n}.$
Appendix C Non-Asymptotic CLT for TSA-PR
We establish finite-time bounds on the Wasserstein-1 distance defined in Eq. (14) between the deviation of the Polyak-Ruppert average $(\bar{x}_{n},\bar{y}_{n})$ around the solution (zeros) and the limiting Gaussian vector. Here we restate Theorem 3.2 for reference.
Theorem C.1.
Let $\bar{z}_{n}\coloneqq(\bar{x}_{n},\bar{y}_{n})$ be the Ruppert-Polyak average of $(x_{t},y_{t})$ generated by (5). Define the asymptotic covariance of $\sqrt{n}(\bar{x}_{n}-x^{*})$ and $\sqrt{n}(\bar{y}_{n}-y^{*})$ as
$\bar{\Sigma}_{ff}=\lim_{n\to\infty}n\mathbb{E}(\bar{x}_{n}-x^{*})(\bar{x}_{n}-x^{*})^{T},\quad\bar{\Sigma}_{ss}=\lim_{n\to\infty}n\mathbb{E}(\bar{y}_{n}-y^{*})(\bar{y}_{n}-y^{*})^{T}.$ (61)
Under Assumptions 13, the above limits exist, and the rate of convergence in the Wasserstein-1 distance is given by:
$\begin{split}d_{1}\left(\sqrt{n}G(\bar{x}_{n}-x^{*}),(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(\frac{n^{a-1/2}}{a-1/2}+\frac{n^{(b-a)/2}}{b-a}+n^{b/2}\right)\right),\\ d_{1}\left(\sqrt{n}\Delta(\bar{y}_{n}-y^{*}),(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(\frac{n^{a-1/2}}{a-1/2}+\frac{n^{(b-a)/2}}{b-a}+n^{b/2}\right)\right),\end{split}$ (62)
where $Z_{1}$ and $Z_{2}$ are standard Gaussian vectors of appropriate dimensions.
The asymptotic covariances $\bar{\Sigma}_{ff}$ and $\bar{\Sigma}_{ss}$ satisfy
$\begin{split}G\bar{\Sigma}_{ff}G^{T}&=\Gamma_{ff}+A_{fs}A_{ss}^{-1}\Gamma_{ss}(A_{fs}A_{ss}^{-1})^{T}-\left(\Gamma_{fs}(A_{fs}A_{ss}^{-1})^{T}+A_{fs}A_{ss}^{-1}\Gamma_{sf}\right)\\ \Delta\bar{\Sigma}_{ss}\Delta^{T}&=\Gamma_{ss}+A_{sf}A_{ff}^{-1}\Gamma_{ff}(A_{sf}A_{ff}^{-1})^{T}-\left(\Gamma_{sf}(A_{sf}A_{ff}^{-1})^{T}+A_{sf}A_{ff}^{-1}\Gamma_{fs}\right)\end{split}.$ (63)
Equivalently,
$\lim_{n\to\infty}\mathrm{Cov}G\bar{x}_{n}=\mathrm{Cov}(W_{1}-A_{fs}A_{ss}^{-1}V_{1}),\quad\lim_{n\to\infty}\mathrm{Cov}\Delta\bar{y}_{n}=\mathrm{Cov}(V_{1}-A_{sf}A_{ff}^{-1}W_{1}).$ (64)
C.1 Proof of Theorem 3.2
Step 1: We write the recursions (5) in the form
$Gx_{t}=(W_{t}-A_{fs}A_{ss}^{-1}V_{t})+\alpha_{t}^{-1}(x_{t}-x_{t+1})-\gamma_{t}^{-1}A_{fs}A_{ss}^{-1}(y_{t}-y_{t+1})$ (65)
$\Delta y_{t}=(V_{t}-A_{sf}A_{ff}^{-1}W_{t})+\gamma_{t}^{-1}(y_{t}-y_{t+1})-A_{sf}A_{ff}^{-1}\alpha_{t}^{-1}(x_{t+1}-x_{t}),$
where $G,\Delta$ are the Schur complements
$G=A_{ff}-A_{fs}A_{ss}^{-1}A_{sf},\quad\Delta=A_{ss}-A_{sf}A_{ff}^{-1}A_{fs}.$
Proof C.2.
Let us start with the fast-time-scale variable update
$x_{t+1}$ $=x_{t}-\alpha_{t}(A_{ff}x_{t}+A_{fs}y_{t}-W_{t})$
$\Rightarrow x_{t}$ $=(\alpha_{t}A_{ff})^{-1}(x_{t+1}-x_{t})-A_{ff}^{-1}(A_{fs}y_{t}-W_{t}).$
Substituting into the slow-time-scale variable’s recursion,
$y_{t+1}$ $=y_{t}-\gamma_{t}\left(A_{sf}x_{t}+A_{ss}y_{t}-V_{t}\right)$
$=y_{t}-\gamma_{t}A_{sf}\left(A_{ff}^{-1}\alpha_{t}^{-1}(x_{t+1}-x_{t})\right)+\gamma_{t}A_{sf}A_{ff}^{-1}A_{fs}y_{t}-\gamma_{t}A_{sf}A_{ff}^{-1}W_{t}-\gamma_{t}A_{ss}y_{t}+\gamma_{t}V_{t}.$
Using $\Delta=A_{ss}-A_{sf}A_{ff}^{-1}A_{fs}$, we have
$y_{t+1}$ $=(I-\gamma_{t}\Delta)y_{t}-\frac{\gamma_{t}}{\alpha_{t}}A_{sf}A_{ff}^{-1}(x_{t+1}-x_{t})+\gamma_{t}(V_{t}-A_{sf}A_{ff}^{-1}W_{t})$ (66)
$\Leftrightarrow\gamma_{t}\Delta y_{t}$ $=y_{t}-y_{t+1}-\frac{\gamma_{t}}{\alpha_{t}}A_{sf}A_{ff}^{-1}(x_{t+1}-x_{t})+\gamma_{t}(V_{t}-A_{sf}A_{ff}^{-1}W_{t}).$
Dividing both sides by the step size $\gamma_{t}$, we have
$\Delta y_{t}=\gamma_{t}^{-1}(y_{t}-y_{t+1})-\alpha_{t}^{-1}A_{sf}A_{ff}^{-1}(x_{t+1}-x_{t})+(V_{t}-A_{sf}A_{ff}^{-1}W_{t}).$ (67)
We repeat the same steps for the fast iterate: Using that
$A_{ss}y_{t}=\gamma_{t}^{-1}(y_{t}-y_{t+1})-(A_{sf}x_{t}-V_{t}),$
we have by substitution
$x_{t+1}$ $=x_{t}-\alpha_{t}(A_{ff}x_{t}-W_{t})-\alpha_{t}A_{fs}y_{t}$
$=x_{t}-\alpha_{t}A_{ff}x_{t}+\alpha_{t}W_{t}-\alpha_{t}A_{fs}A_{ss}^{-1}\left(\gamma_{t}^{-1}(y_{t}-y_{t+1})-(A_{sf}x_{t}-V_{t})\right)$
$=x_{t}-\alpha_{t}\left(A_{ff}-A_{fs}A_{ss}^{-1}A_{sf}\right)x_{t}+\alpha_{t}(W_{t}-A_{fs}A_{ss}^{-1}V_{t})-\frac{\alpha_{t}}{\gamma_{t}}A_{fs}A_{ss}^{-1}(y_{t}-y_{t+1}).$
Denoting $G=A_{ff}-A_{fs}A_{ss}^{-1}A_{sf}$, we have
$Gx_{t}=\alpha_{t}^{-1}(x_{t}-x_{t+1})+(W_{t}-A_{fs}A_{ss}^{-1}V_{t})-\gamma_{t}^{-1}A_{fs}A_{ss}^{-1}(y_{t}-y_{t+1}).$ (68)
The Polyak-Ruppert averages $\bar{x}_{n},\bar{y}_{t}$ are obtained by taking the average on both sides:
$G\bar{x}_{n}$ $=\frac{1}{n}\sum_{t=1}^{n}(W_{t}-A_{fs}A_{ss}^{-1}V_{t})+\frac{1}{n}\sum_{t=1}^{n}\alpha_{t}^{-1}(x_{t}-x_{t+1})-\frac{1}{n}\sum_{t=1}^{n}\gamma_{t}^{-1}A_{fs}A_{ss}^{-1}(y_{t}-y_{t+1})$
$\Delta\bar{y}_{n}$ $=\frac{1}{n}\sum_{t=1}^{n}(V_{t}-A_{sf}A_{ff}^{-1}W_{t})+\frac{1}{n}\sum_{t=1}^{n}A_{sf}A_{ff}^{-1}\alpha_{t}^{-1}(x_{t}-x_{t+1})+\frac{1}{n}\sum_{t=1}^{n}\gamma_{t}^{-1}(y_{t}-y_{t+1}).$
Step 2: Quantitative bounds on the central limit theorem.
Asymptotic convergence of $\sqrt{n}(\bar{x}_{n},\bar{y}_{n})$ to a normal distribution can then be proved by showing that (1) the first noise terms above satisfy Lindeberg’s condition and that (2) the remaining terms converge to zero with probability 1. Slutsky’s theorem then implies that the sum $N_{t}+E_{t}\to N+c$ when the noise $N_{t}$ converges in distribution to a random variable $N$ and the error terms $E_{t}$ converge to a constant $c$ with probability 1.
For non-asymptotic analysis, we use the mean-square results for the last iterates to obtain finite-time bounds on the telescoped weighted differences above, where they converge in probability to 0. Next, we use a bound on the Wasserstein-1 distance between the first terms and a normal random variable. These results can be combined using Lindeberg’s decomposition to obtain finite-time bounds on the distance between $\bar{x}_{n}$ and $\bar{y}_{n}$ to a Gaussian random variable.
Lemma 5.2 resembles Slutsky’s theorem for a sum of two convergent random sequences, but can be used to obtain finite-time bounds on the Wasserstein-1 distance when the random sequences being summed converge with respect to the Wasserstein-1 distance and mean absolute error.
Lemma
(Restated Lemma 5.2) Consider two random sequences $\{X_{t}\},\{Y_{t}\}$ such that $d_{1}(X_{t},X)\leq r_{t}$ for some random variable $X$ and $\mathbb{E}\lVert Y_{t}\rVert\leq r^{\prime}_{t}$. Then, $d_{1}(X_{t}+Y_{t},X)\leq r_{t}+r^{\prime}_{t}$.
Proof C.3.
By the definition (14), we have that
$d_{1}(X_{t}+Y_{t},X)$ $=\sup_{h\in\mathrm{Lip}_{1}}\mathbb{E}[h(X_{t}+Y_{t})-h(X)]$
$=\mathbb{E}\left[h(X_{t}+Y_{t})-h(X_{t})\right]+\mathbb{E}h(X_{t})-\mathbb{E}h(X)$
$\leq\mathbb{E}\lVert Y_{t}\rVert+d_{1}(X_{t},X),$
where the last inequality uses that $h$ in 1-Lipschitz. This statement demonstrates that the Wasserstein-1 distance between a sum and its limit can be decomposed into mean absolute errors and the Wasserstein-1 distance of a summand.
Next, we use a non-asymptotic central limit theorem (Theorem 1, Srikant (2024)) for martingale differences, simplified for our application.
Lemma
(Restated Lemma 5.1) Let $\{N_{t}\}$ be a martingale difference sequence satisfying Assumption 2. Denoting $Z\sim\mathcal{N}(0,I)$ to be the standard Gaussian vector, we have that
$d_{1}\left(n^{-1/2}\sum_{t=1}^{n}N_{t},\Gamma^{1/2}Z\right)\leq\mathcal{O}\left(\frac{d}{1-\beta}\right)\frac{\lVert\Gamma^{1/2}\rVert}{n^{\beta/2}}\left(\lVert\Gamma^{-1/2}\rVert^{2+\beta}+\lVert\Gamma^{-1/2}\rVert^{\beta}\right).$ (69)
It remains to show that the expected norm of the remaining terms decay to zero. Observe that
$\sum_{t=1}^{n}\alpha_{t}^{-1}(x_{t}-x_{t+1})=\frac{x_{1}}{\alpha_{1}}-\frac{x_{n+1}}{\alpha_{n}}+\sum_{t=1}^{n-1}(\alpha_{t+1}^{-1}-\alpha_{t}^{-1})x_{t}.$
The step sizes in Assumption 3 satisfy
$\alpha_{t+1}^{-1}-\alpha_{t}^{-1}\leq(t\alpha_{t})^{-1},$
and we have from triangle inequality that
$\mathbb{E}\lVert\sum_{t=1}^{n}\alpha_{t}^{-1}(x_{t}-x_{t+1})\rVert\leq\alpha_{1}^{-1}\mathbb{E}\lVert x_{1}\rVert+\alpha_{n}^{-1}\mathbb{E}\lVert x_{n+1}\rVert+(\alpha_{1})^{-1}\sum_{t=1}^{n}t^{a-1}\mathbb{E}\lVert x_{t}\rVert.$ (70)
Repeating for $\{y_{t}\}$, we have
$\mathbb{E}\lVert\sum_{t=1}^{n}\gamma_{t}^{-1}(y_{t}-y_{t+1})\rVert\leq\gamma_{1}^{-1}\mathbb{E}\lVert y_{1}\rVert+\gamma_{n}^{-1}\mathbb{E}\lVert y_{n+1}\rVert+(\gamma_{1})^{-1}\sum_{t=1}^{n}t^{b-1}\mathbb{E}\lVert y_{t}\rVert.$ (71)
Next, we obtain bounds on $\mathbb{E}\lVert x_{t}\rVert$ and $\mathbb{E}\lVert y_{t}\rVert$ from Theorem 3.1 and Jensen’s inequality. Following the proof of Theorem 3.1, the dominant rates that appear in the error bounds for TSA (without PR averaging) are given by
$\lVert\mathbb{E}(x_{t+1}-x_{\infty}(y_{t+1})+L_{t}\hat{y}_{t+1})(x_{t+1}-x_{\infty}(y_{t+1})+L_{t}\hat{y}_{t+1})^{T}-\alpha_{t+1}\Sigma_{ff}\rVert=\mathcal{O}\left(\gamma_{t}+\frac{1}{t}\right)$
$\lVert\mathbb{E}(x_{t+1}-x_{\infty}(y_{t+1})+L_{t}\hat{y}_{t+1})\hat{y}_{t+1}^{T}-\gamma_{t+1}\Sigma_{fs}\rVert=\mathcal{O}\left(\alpha_{t}^{-1}\left(\frac{\gamma_{t}}{t}+\gamma_{t}^{2}+\alpha_{t}^{2}\gamma_{t}\right)+\frac{\gamma_{t}}{t}+\gamma_{t}^{2}\right)$
$\lVert\mathbb{E}\hat{y}_{t+1}\hat{y}_{t+1}^{T}-\gamma_{t+1}\Sigma_{ss}\rVert=\mathcal{O}\left(\frac{1}{t}+\alpha_{t}\gamma_{t}\right).$
We are interested in obtaining $\mathbb{E}\lVert x_{t}\rVert^{2}$ and $\mathbb{E}\lVert y_{t}\rVert^{2}$ when $x^{*},y^{*}$ are zero (due to zero bias assumed for simplicity). Hence, $\hat{y}_{t}=y_{t}$ and we have the square norm of the iterates
$\lVert\mathbb{E}y_{t+1}y_{t+1}^{T}\rVert$ $=\mathcal{O}\left(\gamma_{t}+\frac{1}{t}+\alpha_{t}\gamma_{t}\right)$
$\lVert\mathbb{E}x_{t+1}x_{t+1}^{T}\rVert$ $=\mathcal{O}\left(\alpha_{t}+\gamma_{t}+\frac{1}{t}+\lVert\mathbb{E}x_{t+1}y_{t+1}\rVert+\mathbb{E}\lVert y_{t+1}y_{t+1}^{T}\rVert\right).$
Combining, it holds that
$\mathrm{Tr}\mathbb{E}x_{t+1}x_{t+1}^{T}$ $\leq\mathrm{dim}(x)\lVert\mathbb{E}x_{t+1}x_{t+1}^{T}\rVert=\mathcal{O}\left(\alpha_{t}+\gamma_{t}+\frac{1}{t}+\alpha_{t}\gamma_{t}\right),$
$\mathrm{Tr}\mathbb{E}y_{t+1}y_{t+1}^{T}$ $\leq\mathrm{dim}(y)\mathbb{E}\lVert y_{t+1}y_{t+1}^{T}\rVert=\mathcal{O}\left(\gamma_{t}+\frac{1}{t}+\alpha_{t}\gamma_{t}\right)$
excluding some non-dominant terms when $b=a+1/\log n$ and $a=1/2+1/\log n$ for large $n$. Using $\mathbb{E}\lVert x_{t}\rVert\leq\sqrt{\mathrm{Tr}\mathbb{E}x_{t}x_{t}^{T}}$, we obtain bounds on $\mathbb{E}\lVert x_{t}\rVert$ and $\mathbb{E}\lVert y_{t}\rVert$.
To complete the error bound, we substitute the above into Eqs. (70) and (71) and simplify as
$\frac{1}{\sqrt{n}}\sum_{t=1}^{n}t^{a-1}\mathbb{E}\lVert x_{t}\rVert=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(n^{a/2}+\frac{n^{a-1/2}}{a-1/2}+\frac{1}{b-a}\right)\right),$
$\frac{1}{\sqrt{n}}\sum_{t=1}^{n}t^{b-1}\mathbb{E}\lVert y_{t}\rVert=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(n^{b/2}+\frac{n^{b-1/2}}{b-1/2}+\frac{n^{(b-a)/2}}{b-a}\right)\right).$
Combining, $d_{1}\left(\sqrt{n}G\bar{x}_{n},(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\right)$ and $d_{2}\left(\sqrt{n}(\Delta\bar{y}_{n},(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/}Z_{2}\right)$ are dominated by the sum of the above terms which is simplified as
$\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(\frac{n^{a-1/2}}{a-1/2}+\frac{n^{(b-a)/2}}{b-a}+n^{b/2}\right)\right).$
Therefore, with the choice of step sizes $a=1/2+1/\log n$ and $b=a+1/\log n$ when $n$ is sufficiently large, we obtain the error bound $\mathcal{O}(n^{-1/4})$.